### v1

# Reducing Vision Transformer Latency on Edge Devices via GPU Tail Effect and Training-free Token Pruning

Abstract
This paper investigates how to efficiently deploy transformer-based neural networks on edge devices. Recent methods reduce the latency of transformer neural networks by removing or merging tokens, with small accuracy degradation. However, these methods are not designed with edge device deployment in mind, and do not leverage information about the hardware characteristics to improve efficiency. First, we show that the relationship between latency and workload size is governed by the GPU tail-effect. This relationship is used to create a token pruning schedule tailored for a pre-trained model and device pair. Second, we demonstrate a training-free token pruning method utilizing this relationship. This method achieves accuracy-latency trade-offs in a hardware aware manner. We show that for single batch inference, other methods may actually increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%. For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
1 Introduction
In the past decade, Internet of Things (IoT) and commodity edge devices for various applications have become ubiquitous [4, 19, 38]. Edge devices have become sufficiently powerful, and model miniaturization techniques sufficiently capable, that machine learning models can be deployed to the network edge for various tasks, including computer vision applications [11, 32]. However, state of the art performance on various computer vision tasks are often claimed by Vision Transformer (ViT) based neural network [10] architectures. ViT models such as DeiT [34] or backbones such as DINOv2 [25] are not designed nor miniaturized for edge deployment.
Prior work [20] has shown that ViT’s feature high levels of redundancy that can be exploited for latency reduction benefits. One general approach is to identify parts of the input (tokens) which contain low information content and remove them. This is referred to as token sparsification. Training-free token sparsification methods such as Token Merging (ToMe) [1] have been effective at reducing latency of pre-trained models. Other approaches like DynamicViT [30] can yield better accuracy than training-free methods, but at the cost of training time on server-grade hardware.
Figure 1: Forward pass latency for widely used DeiT-B and DinoV2 models across various hardware (See Tab. 3) evaluated on the ImageNet1K [8] dataset. These plots demonstrate the relationship between workload size and latency can be variable and non-linear across a variety of hardware. Consequently, it is possible to achieve large latency reductions without removing too many tokens in many cases. Note that the latency axis is not necessarily indexed from 0 for all plots.
We address three shortcomings of existing techniques. (1) Many existing efficient methods do not consider fine-grained hardware or latency characteristics [1, 16, 37]. Fig. 1 demonstrates the diversity of latency-workload trends across devices and workload sizes. As a result, there is room to optimize efficient methods in a hardware aware manner by considering this relationship. (2) Some existing efficient methods may require extensive training [42, 30, 5] hinders the deployment of pre-trained models on edge devices. (3) Prior work has investigated hardware aware methods for CNN’s [41, 40], the GPU tail effect has not been investigated specifically utilized for transformer models. Given these properties of existing token removal methods, this paper focuses on reducing ViT latency by considering hardware characteristics and without requiring training. This paper has the following contributions:
  1. 1.
    Introduce the first method for deciding a hardware-aware ViT token pruning schedule for model and device pairs.
  2. 2.
    Develop a novel training-free token pruning mechanism. For similar latency (within 5.2%) across various hardware and workload sizes, we achieve 0.46-43.7 higher accuracy than ToMe, a popular state-of-the-art method.
2 Background and Related Work
In this section, we review related work on model acceleration and efficient methods for Vision Transformer (ViT) models. Some post-processing or fine-tuning methods such as quantization [18] are compatible with token sparsification techniques such as our method.
2.1 Model Acceleration of Vision Transformers
ViT  [10] architectures such as DINOv2 [25] have achieved state-of-the-art accuracy on multiple computer vision tasks, including image classification and object detection. State-of-the-art models such as DINOv2-G [25] have over 1 billion parameters. It is important to address the efficiency of these models when deploying on edge devices.
There are many existing model acceleration techniques for ViT’s such as quantization [18], knowledge distillation [39], low-rank factorization [6], and optimized kernels [7] for attention operations. These methods often aim to reduce the complexity of the transformer attention mechanism, or learn a compressed representation that has similar effectiveness to a higher-performing model or representation.
One approach to reduce the latency of transformer models is token sparsification, removing less-important tokens [1, 15, 31, 30] to reduce workload size. These methods are easily applied to a variety of ViT architectures [34, 25, 10]. A potential benefit of sparsification approaches is that they do not necessarily require any training or fine-tuning [36, 1, 2] . Some methods [30, 42, 5, 37] may require significant training time, e.g. 100+ epochs, to recover more accuracy and yield latency reductions. Training poses a high barrier to application; training-free methods are more accessible [1, 36].
2.2 The GPU Tail Effect
The relationship between latency and workload size can be non-linear as seen in Fig. 1. Large changes in latency are caused by the GPU Tail Effect [40], arising due to workload partitioning and dispatching to the cores of a device’s GPU [23]. If, at any layer, the GPU cannot process the entire workload at once another wave of computation will be launched. A partially full wave costs the same as a full wave. The overhead introduced by a partial wave accumulates on a layer-wise basis. Consequently, a minor change in workload size may lead to disproportionately large change in latency since the overhead of partial waves is reduced across all layers. For example, on the NVIDIA AGX Orin, removing one token can change latency by up to 33% (Fig. 1).
Prior work has exploited the tail effect to guide efficient methods for machine learning models [41, 40], particularly in the context of convolutional neural network (CNN) channel pruning [13]. These methods often reduce latency with negligible accuracy degradation. However, ViT token pruning mechanisms are fundamentally different from CNN channel pruning approaches due to architectural differences between CNN’s and ViT’s. Therefore, we hypothesize that the GPU tail effect can lower latency for ViT’s, and recognize that new techniques will be needed. Our method validates this hypothesis.
3 Token Pruning with Consideration of the GPU Tail Effect
In Sec. 2.1 and Sec. 2.2, we discussed the convenience factor of training-free token pruning methods and how previous work leveraged the GPU tail effect for efficiency benefits. Therefore, we choose two design goals for our method: (1) require no training or fine-tuning of a pre-trained ViT model; and (2) achieve more and/or better accuracy-latency tradeoffs by pruning tokens according to the GPU tail effect. Tab. 1 illustrates qualitative differences between our work and others as a result of these goals.
Method Training Free? Hardware Aware?
Ours
ToMe [1]
Zero-TPrune [36]
Top-K [12],[16]
DynamicViT [30]
STViT [5]
TPS [37]
Table 1: Qualitative differences between our pruning approach and similar techniques.
As depicted in Fig. 2, our token pruning approach consists of two main parts. First, we determine a pruning schedule for a particular model and device pair — this includes identifying the number of tokens to remove and the layers at which to remove them. Second, we develop a training-free pruning technique to remove non-informative tokens during inference time. In Sec. 3.1 we show how to decide the number of tokens to prune based on the GPU tail effect. Next, in Sec. 3.2 we explain our choice of which layers to prune, completing our offline pruning schedule selection. Finally, Sec. 3.3 describes our token pruning mechanism which is used at inference time.
Figure 2: Illustration of our method to decide a pruning schedule (left) and how we prune according to the schedule at inference time (right).
3.1 Deciding a Number of Tokens to Prune for our Pruning Schedule
Workload size definition. Before explaining how to leverage the GPU tail effect, we first define workload size. We model ViT workload size as being proportional to the number of input tokens $n$, embedding dimension $d$, and batch size $b$ in Equation (1). This is because the runtime complexity of the self-attention mechanism is $O(n^{2}\times d)$ [35]; thus the workload size is proportional to $n$ and $d$. The batch size $b$ is simply a multiplier for workload size. Transformer models are (partially) parameterized by a particular choice of $d$. For DeiT-S, $d$ is $384$; for DinoV2-G, $d$ is $1536$. Fig. 1 illustrates the workload size vs latency relationships by varying $n$, $d$, and $b$ across a set of devices. Eq. 1 describes the relationship in terms of the same quantities:
$\text{Workload Size}\propto n\times d\times b.$ (1)
We measure and leverage the GPU tail effect’s characteristics for a particular workload size and device to expand the set of (better) accuracy-latency tradeoffs for token pruning.
Measuring the GPU tail effect rather than predicting it. Prior work has modeled GPU tail effect behavior for a convolutional neural network with some assumptions [17]: (1) The latency staircase pattern increases monotonically; (2) The latency difference between steps is uniform. Our experiments indicate that these assumptions may not hold for certain workload sizes using transformer networks.
First, model latency may actually increase as workload size decreases for certain workload sizes. We demonstrate this later in Sec. 4. Fig. 1 (a) demonstrates this phenomenon for DeiT-B on the TX2. This anomaly was due to PyTorch [27] launching a different set of kernels that had higher latency for one set of workload sizes.
Second, the latency difference between steps can be nonuniform. The distance between steps varies across workload sizes as seen in Fig. 1 (a-c). Considering these observations, we empirically measure the GPU tail effect due to the difficulty in predicting or modeling its behavior. The next part describes the necessary measurements.
Assigning utility to tokens for deciding a pruning schedule. Given a pre-trained model $M$ with a number of input tokens $N$, we frame selection of the number of tokens to prune $R$ as a multi-objective optimization problem between latency gains and accuracy degradation. This is the offline computation related to selecting our pruning schedule for $M$ on a target device, as seen in Fig. 2.
First, we measure latency $L(n)$ and accuracy degradation $A(n)$ of $M$ on the target device for each number of tokens to keep $n\in[1,2,...,N]$. Measuring latency of $M$ is performed with a grid search across $n$. For estimating accuracy degradation, we need a heuristic that does not depend on our pruning schedule selection. Furthermore it is useful to underestimate the accuracy of $M$ so our selection algorithm is hesitant to remove too many tokens, which can degrade accuracy significantly. A simple proxy for the accuracy of each $n$ is to apply random token removal after the first layer on $M$ [26]. It is assumed that any token pruning method should be better than random token removal since random token pruning does not consider token information content at all. Furthermore, pruning at the first layer will degrade accuracy more than pruning later in the network. Therefore random token pruning is a suitable heuristic for estimating accuracy degradation.
[…]
$\text{U}_{\text{A}}(n)=\frac{A(n)}{\text{max}\,A(n)}$ (2)
$\text{U}_{\text{L}}(n)=1-\frac{L(n)}{\text{max}\,L(n)}$ (3)
[…]
Fig. 3 illustrates an example of the plots of $L(n)$, $A(n)$ and the overall utility $U(n)$. With $N=257$, we follow Eq. 4 to obtain $N_{\text{keep}}=91$ and $R=257-91=166$. As part of our evaluation, we measure the effect of different values of $\alpha$ (Sec. 4.2).
Figure 3: Example offline computation to decide the number of tokens prune. We find $n=N_{\text{keep}}$ with maximum utility and by extension the number of tokens to prune $R$ per Eq. 4. Configuration: $\alpha$=0.5, $M$=DinoV2-G, and AGX Orin.
3.2 Pruning Schedule: Deciding Layers at which to Prune
Now we explain the layers to which our pruning mechanism will be applied. This is the final step of the offline computation illustrated in Fig. 2.
Our schedule prunes all $R$ tokens at one layer, fairly early in the network. This differs from other methods such as ToMe [1], Top-K [12, 16], and DynamicViT [30] that progressively prune tokens from layer to layer. Our design is based on two observations supported by our evaluation. First, on small, edge-grade workloads, repeated pruning operations themselves can introduce non-trivial latency (Sec. 4.5). Second, the GPU Tail Effect will lead to improved latency for each layer after which pruning has occurred — earlier pruning means more layers will benefit (Sec. 4.3).
Fig. 4 illustrates the difference between single-layer and progressive pruning. Our approach attempts to place the token pruning at or near the first 25% of layers, similar to the first pruning layer of DynamicViT [30]. For DinoV2-G, we prune at the 10th layer, and for DeiT models, we prune at the 4th layer. This allows us to realize latency reductions according to the GPU tail effect for the remaining 30 layers in DinoV2-G and the 8 remaining layers for DeiT-S and DeiT-B. We discuss the effect of pruning location in Sec. 4.3.
Figure 4: Illustration of the difference between our pruning mechanism and Token Merging [1] ToMe increases latency of DeiT-S for single batch inference on a Jetson TX2 [24] due to the overhead of pruning at every layer. Our method is able to yield a latency reduction by pruning fewer tokens at once earlier in the network.
3.3 Token Pruning Method
The schedule selection described in Secs. 3.1 and 3.2 identifies the location and number of tokens to prune. Now we give a training-free token pruning mechanism to decide which tokens to prune. This pruning mechanism is applied during inference (Fig. 2). The pruning schedule, computed offline, is an input to our token pruning mechanism.
In terms of design goals, we want our pruning mechanism to require (1) no training or finetuning on $M$, and (2) to introduce minimal overhead. First, requiring no fine-tuning or training is a convenience for practitioners, and allows pre-trained models to be deployed readily. Second, keeping overhead to a minimum for small workload sizes is necessary to yield latency reductions as previously mentioned in 3.2.
The primary design choice of pruning techniques is choosing a method to rank the importance of tokens. Similarly to other pruning techniques [3, 14], we rank the importance of each token $i$ by measuring the quantity of attention received from all other tokens. This is analogous to summing along the columns of the attention matrix in a transformer encoder layer.
Our method incorporates two aspects of importance, one from the $A_{\text{H}}$ matrix and one from the $V$ matrix. For attention, we proceed as follows. Recall that the core attention mechanism of transformers [35] is parameterized by the $Q$, $K$, and $V$ matrices. The attention matrix (including softmax) of a transformer encoder layer with $N$ tokens and $H$ heads is given by $A_{\text{H}}\in\mathbb{R}^{H\times N\times N}$  [35]. To convert the softmax attention matrix, $A_{\text{H}}$, into tokenwise importance scores we apply a reduction along the along the head dimension, $H$. We find the maximum along the head dimension for $A_{\text{H}}$. We found that applying max marginally increased accuracy compared to mean or sum. This reduction yields a new matrix $A\in\mathbb{R}^{N\times N}$. Now, summing across rows of $A$ yields the quantity of attention each token $i$ receives, similar to a component of EViT’s pruning mechanism [16] or Top-K [12]. This yields an importance score $\text{A}_{\text{S}}[i]\in\mathbb{R}^{N}$ for each token $i$.
$V_{\text{S}}$ Feature Choice Feature ImageNet1K Accuracy (%) Q 79.74 K 79.77 V 79.81
$V_{\text{S}}$ Head Reduction Operator Feature ImageNet1K Accuracy (%) mean 79.72 sum 79.70 max 79.81
Table 2: Token pruning design ablations using for our $V_{S}$ term our schedule on DeiT-S-dist for batch-size 2. In our code we use torch.amax [29] instead of max to handle the batch dimension.
Figure 5: Illustration of our token pruning method for a single transformer encoder layer at inference time. We compute an attention score and V score for each token to estimate importance, then prune the least informative tokens. Pruned tokens are aggregated into a single ‘garbage bin’ or inattentive token to retain information content as in EViT [16].
We then consider a second importance term based on the $V\in\mathbb{R}^{H\times N\times C}$ matrix. This term assigns higher scores to tokens whose entry in the $V$ matrix has a large magnitude compared with other tokens. Adding this term can marginally improve accuracy. Similarly to the attention matrix, we find the maximum along the $H$ dimension [29] — we do a set of ablations summarized in Tab. 2. Post reduction, we have $V^{\prime}\in\mathbb{R}^{N\times C}$ and we have:
$\text{V}_{\text{S}}[i]$ $=\text{softmax}(\sum^{C}_{c=1}V^{\prime}[i,c])\in\mathbb{R}^{N}.$ (5)
We add both of these scores to yield our importance score for each token $i$:
$\text{Importance}[i]$ $=\text{A}_{\text{S}}[i]+\text{V}_{\text{S}}[i].$ (6)
Finally, we borrow from EViT [16] and instead of discarding the pruned tokens, we create a new “inattentive” token based on the averaged features of all pruned tokens and append it to the set of kept tokens. This allows information from pruned tokens to be preserved in this inattentive token, which EViT found improved accuracy. Fig. 5 illustrates the token pruning process for a ViT encoder block or layer. Note that we do not prune specialized tokens such as the CLS [9] or DIST [33] tokens. These tokens serve additional purposes beyond representing the input.
4 Evaluation
In this section, we demonstrate the effectiveness of our token pruning schedule and pruning mechanism across devices (listed in Tab. 3) and across various workload sizes. To justify our design decisions, we investigate our $\alpha$ hyperparameter and pruning location choice for our schedule selection and pruning mechanism in Sec. 4.2 and Sec. 4.3, respectively. Then, we compare with a variety of existing token removal methods such as DynamicViT [30], ToMe [1]. We use Top-K [16, 12] as a baseline on ImageNet1K [8]. Sec. 4.4 demonstrates how our method advances the accuracy vs. latency pareto front across devices, workload sizes, and models. Finally, we identify some pitfalls of existing token pruning techniques for small workload sizes in Sec. 4.5. Particularly we show that some existing methods actually increase latency rather than reduce it.
4.1 Experimental Setup
Evaluation is done using three devices with varying hardware characteristics (Tab. 3). These devices feature various generations of NVIDIA GPU architectures, and thus feature different levels of computing power. We use NVIDIA TX2 [24] and NVIDIA AGX Orin [21] development boards which are edge devices designed for machine learning. Additionally, evaluation is done on a server-grade system with an NVIDIA A100 GPU [22] to demonstrate our method can generalize across a wide range of hardware, even though we design our method for edge devices. Experiments performed on the TX2 and AGX Orin were performed with fixed CPU and GPU clock rates for reproducibility and consistency. The A100 system clocks were unable to be locked because the servers machines are university-wide resources.
Device RAM (GB) GPU Cores CUDA Cores Max Power (W)
TX2 [24] 8 2 256 15
AGX Orin [21] 32 14 1792 40
A100 [22]  ✝ 80 108 6192 300
Table 3: Summary of device hardware information. We evaluate our method on two edge devices (TX2 and AGX Orin) and one server-grade system in this work.  ✝ The A100 power consumption is the power consumption only for the GPU, not the entire system.
In order to measure latency for evaluation, we use the PyTorch benchmark module [28]. All accuracy measurements were performed on the A100 system using a batch size of 512. Conveniently, accuracy only needs to be measured once for a particular pre-trained model $M$. It is only necessary to record the latency $L(n)$ on a target device for a particular workload size in order to capture GPU tail effect behavior. In subsequent experiments, we decide the pruning hyperparameters of each token pruning method for fair comparison with our method.
For reproducibility, our code is open source. We include our method, and modifications to the original ToMe codebase in order to support DinoV2 pre-trained models.
4.2 Pruning schedule ablation study across $\alpha$ (utility weight of latency/accuracy)
In 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect. The hyperparameter of this algorithm is $\alpha$, which determines the relative weighting of the accuracy and latency utilities. Setting to $\alpha=1$ prioritizes accuracy, and only a few tokens might be pruned. Setting to $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens. To understand the performance spectrum, we perform an ablation study. The data is in Tab. 4 and illustrated in Fig. 6.
Device $\alpha$ Range Computed $R$ Accuracy Median
(# tok. to remove) Degradation Latency (ms)
AGX Orin 0.1-0.3 139 1.07 112.2 (-27.8%)
0.4-0.5 166 2.04 104.2 (-33.0%)
0.6-0.9 193 4.82 100.9 (-35.1%)
A100 0.1-0.4 73 0.25 28.77 (-4.70%)
0.5-0.9 139 1.07 28.27 (-6.34%)
Table 4: A companion table for Fig. 6 that shows the accuracy vs latency tradeoffs for the computed number of tokens to prune $R$ for various $\alpha$. Latency percent change is computed with respect to the unmodified DinoV2-G inference time.
Figure 6: Computed pruning rate $R$ from our offline pruning schedule versus various $\alpha$. Note that the selection of $R$ is relatively invariant to changes in $\alpha$. Configuration: $M$ = DinoV2-G and batch size = 2.
Evaluation on the AGX Orin for $\alpha>0.5$ yielded $R=193$ (over 75% of tokens pruned), which degraded accuracy by $\sim$2.8for 2.1% less latency. We consider this to be an unfavorable tradeoff. This ablation, in addition to results in Sec. 4.4, suggest that $\alpha\leq 0.5$ is a good choice. Thus we use $\alpha=0.5$ for all evaluations that appear in this work. Intuitively speaking this means accuracy degradation and latency reduction are considered with equal weight according to Eq. 4 when selecting $R$.
4.3 Pruning location ablation study
In 3.2 we decide at which layer our pruning mechanism should be applied. To provide insight into the potential pruning locations, we performed an ablation study. Tab. 5 illustrates latency and accuracy tradeoffs for various pruning locations of DinoV2-G.
Batch Size Pruning Accuracy Median
Layer Degradation Latency (ms)
1 1/40 3.19 68.4
10/40 1.07 81.3
20/40 0.58 93.4
30/40 0.49 104.4
2 1/40 7.08 79.1
10/40 2.04 104.7
20/40 0.97 133.0
30/40 0.83 160.8
Table 5: Latency/accuracy tradeoffs by pruning location. Configuration: $M$ = DinoV2-G on AGX Orin.
As expected, pruning earlier yields lower latency but greater accuracy degradation. For the batch size 1, our method pruned $\sim$54of input tokens at the first layer degraded accuracy by 3.19% but yielded a 40% overall latency reduction. Across both batch size 1 and 2 in this ablation study, pruning after the first 25% of layers (layer 10) results in a good balance between latency reduction and accuracy degradation. Pruning later in the network will reduce accuracy degradation, however we prioritize yielding latency benefits with our method. Therefore, we perform pruning at the layer 25% of the way into the network for all models evaluated in this work. That is layer 3 for DeiT-B and DeiT-S, and layer 10 for DinoV2-G.
4.4 Accuracy-latency tradeoffs of our method
In this section, we discuss the results of our training-free token pruning mechanism when combined with the pruning schedule inspired by the GPU tail effect. We primarily compare against ToMe [1], and include Top-K as a baseline [12, 16]. DynamicViT [30] is included in select evaluations, however unlike the other compared work and our method it requires training. Tab. 6 illustrates our method’s ability to retain higher accuracy for similar latency across various devices and workload sizes. Fig. 7 demonstrates that our token pruning mechanism and schedule expands the accuracy-latency tradeoffs on the pareto front across various workload sizes and devices.
First, our method is able to achieve higher accuracy than other training-free pruning techniques ToMe and Top-K. Tab. 6 is an ablation study across various workload sizes (batch size, models) and devices, where we tune the hyperparameters of methods such that similar latency is achieved. Unsurprisingly, DynamicViT is able to retain higher accuracy than all methods, since it requires training.
Device Batch Model and Accuracy Median
Size Method Degradation Latency (ms)
TX2 2 DeiT-B + Top-K 2.44 143.0 (-33.5%)
DeiT-B + ToMe 1.63 142.7 (-33.7%)
DeiT-B + Ours 1.16 146.0 (-32.1%)
DeiT-B + DyViT ✝ 0.57 143.7 (-33.2%)
DeiT-B 215.0 -(00.0%)
TX2 2 DeiT-S + Top-K 4.30 50.32 (-27.0%)
DeiT-S + ToMe 2.47 52.98 (-23.1%)
DeiT-S + Ours 1.24 49.44 (-28.3%)
DeiT-S + DyViT ✝ 0.46 50.88 (-26.2%)
DeiT-S 68.92 -(00.0%)
AGX Orin 2 DinoV2-G + Top-K 45.66 104.4 (-32.9%)
DinoV2-G + ToMe 6.96 104.4 (-32.9%)
DinoV2-G + Ours 2.04 104.1 (-33.1%)
DinoV2-G 155.5 -(00.0%)
A100 4 DinoV2-G + Top-K 13.31 33.76 (-16.7%)
DinoV2-G + ToMe 6.96 33.79 (-16.6%)
DinoV2-G + Ours 2.04 32.37 (-20.1%)
DinoV2-G 40.53 -(00.0%)
Table 6: Comparison with existing methods for similar latency across device and workload sizes. We tune the pruning hyperparameters of the compared methods so they achieve similar latency to our method. ✝DynamicViT (DyViT) requires training. Other evaluated methods are training-free.
Compared with ToMe and Top-K however, it is shown that our method yields consistently lower accuracy degradation. Batch size 2 inference on the AGX Orin demonstrates that to achieve similar latency to our method, Top-K degrades accuracy by more than 45.6%, while we degrade accuracy by only 2%. Across all workload sizes and devices, ToMe featured 0.47%-5.92% lower accuracy than our method (Tab. 6).
Figure 7: Illustration of accuracy and latency tradeoffs of surveyed methods with $M$ = DinoV2-G. Our pruning schedule selection pruning mechanism generate points that expand the pareto front. This indicates that considering the GPU tail effect can provide strictly better tradeoffs than existing methods in some cases. Note that the number of tokens removed at each layer ($r$) of Top-K and ToMe is evaluated from $r=5$ to $=8$ in increments of 1.
Second, we note that for larger pruning rates such as $r=7$ and $r=8$, ToMe and Top-K remove nearly all input tokens by the last layer of DinoV2-G. These high pruning ratios yield high accuracy degradation of over 40% in the case of A100 batch-size 4 for ToMe, as seen in Fig. 7. Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we simultaneously achieve higher accuracy and lower latency than progressive pruning approaches. This shows aggressive pruning at early layers can be strictly better than progressive pruning methods for certain workload sizes. Our token pruning mechanism and schedule offers more accuracy latency tradeoffs across workload sizes.
4.5 Low workload size observations
We demonstrate there are some cases where pruning techniques actually increase latency with respect to the baseline model, rather than reduce it. Tab. 7 demonstrates an example on the TX2 for single batch inference. ToMe increased latency by $\sim$30%, while we reduced latency by $\sim$9%. Our tradeoff of 0.99% accuracy degradation and 9% latency reduction is better than ToMe and Top-K, both of which degrade accuracy and provide no latency benefit compared with the baseline DeiT-S pre-trained model. For certain workload sizes, using a baseline model or a hardware aware method such as ours is strictly better than attempting to use a token removal method with high overhead.
Device Batch Model and Accuracy Median
Size Method Degradation Latency (ms)
TX2 1 DeiT-S + Top-K (r=6) 0.73 43.38 (+18.6%)
DeiT-S + ToMe (r=6) 0.27 50.70 (+30.3%)
DeiT-S + Ours (R=69) 0.99 32.12 (-9.06%)
DeiT-S 35.32 (-0.00%)
Table 7: Given comparable total pruned tokens (69 to 72), for certain workload sizes a token pruning mechanism may increase latency due to the overhead of performing pruning at multiple layers. Our pruning method is applied at a single layer to avoid this relatively significant overhead.
We found that the operations required to do token pruning add a non-trivial latency overhead for small workload sizes on some devices such as the TX2. The overhead from masking or removing tokens can outweigh the latency reduction from processing fewer tokens. As explained in Sec. 3.2, our choice to perform single-layer pruning rather than progressive pruning allows us to yield a latency reduction.
5 Conclusion
In this work, we investigated how to improve the efficiency of pre-trained vision transformers when deployed on edge devices and provided three primary contributions. First, we provided insight into how to determine a token pruning schedule for transformer models by leveraging information about the GPU tail effect. Second, a token pruning approach was introduced that is training-free and has low overhead compared with other pruning methods. We show that our method advances the accuracy-latency pareto front compared with existing token pruning approaches across workload sizes and devices. Furthermore we show that for single batch inference, other methods may actually increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%. For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
6 Limitations
Our approach to token pruning is designed to achieve better tradeoffs at small workload sizes than other methods. This is ideal when inputs are streamed to the model, or if only small batch inference is possible. For larger workload sizes, such as running DinoV2-G with batch sizes of 32 or higher, our method may not be ideal.
7 Future Work
We demonstrate that single layer pruning can be effective even for earlier pruning layers. Learning based approaches may be able to take advantage of the high latency reductions of aggressive, early pruning while maintaining accuracy. Another promising direction is investigating the benefits of composing token sparsification and other acceleration methods together.

### v2

@@
-Reducing Vision Transformer Latency on Edge Devices via GPU Tail Effect and Training-free Token Pruning
+Pruning One More Token is Enough: Leveraging Latency-Workload Non-Linearities for Vision Transformers on the Edge
 Abstract
-This paper investigates how to efficiently deploy transformer-based neural networks on edge devices.
+This paper investigates how to efficiently deploy vision transformers on edge devices.
 Recent methods reduce the latency of transformer neural networks by removing or merging tokens, with small accuracy degradation.
-However, these methods are not designed with edge device deployment in mind, and do not leverage information about the hardware characteristics to improve efficiency.
-First, we show that the relationship between latency and workload size is governed by the GPU tail-effect.
-This relationship is used to create a token pruning schedule tailored for a pre-trained model and device pair.
-Second, we demonstrate a training-free token pruning method utilizing this relationship.
-This method achieves accuracy-latency trade-offs in a hardware aware manner.
-We show that for single batch inference, other methods may actually increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%.
+However, these methods are not designed with edge device deployment in mind: jthey do not leverage information about the latency vs.
+workload trends to improve efficiency.
+First, we show the latency-workload size relationship is nonlinear for certain workload sizes.
+We consider this relationship to create a token pruning schedule.
+Second, we demonstrate a training-free, token pruning method utilizing this schedule.
+We show that for single batch inference, other methods increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%.
 For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
 1 Introduction
-In the past decade, Internet of Things (IoT) and commodity edge devices for various applications have become ubiquitous [4, 19, 38].
-Edge devices have become sufficiently powerful, and model miniaturization techniques sufficiently capable, that machine learning models can be deployed to the network edge for various tasks, including computer vision applications [11, 32].
-However, state of the art performance on various computer vision tasks are often claimed by Vision Transformer (ViT) based neural network [10] architectures.
-ViT models such as DeiT [34] or backbones such as DINOv2 [25] are not designed nor miniaturized for edge deployment.
-Prior work [20] has shown that ViT’s feature high levels of redundancy that can be exploited for latency reduction benefits.
-One general approach is to identify parts of the input (tokens) which contain low information content and remove them.
-This is referred to as token sparsification.
+In the past decade, Internet of Things (IoT) and commodity edge devices have become ubiquitous [4, 22, 43].
+Edge devices have become sufficiently powerful, and model miniaturization techniques sufficiently capable, that machine learning (ML) models can be deployed to the network edge for various tasks, including computer vision applications [12, 38].
+However, state of the art performance on various computer vision tasks are often claimed by large vision transformer (ViT) based neural network [10] architectures.
+ViT models such as DeiT [39] and DINOv2 [30] are not designed nor miniaturized for edge deployment.
+Additionally, latency is often of critical importance on edge devices, and only small inputs or workloads can be processed [18].
+Prior work has shown that ViT’s have high redundancy that can be exploited for latency reduction benefits [23].
+One general approach is to identify and remove parts of the input (tokens) with low information content.
+This is called token sparsification.
 Training-free token sparsification methods such as Token Merging (ToMe) [1] have been effective at reducing latency of pre-trained models.
-Other approaches like DynamicViT [30] can yield better accuracy than training-free methods, but at the cost of training time on server-grade hardware.
-Figure 1: Forward pass latency for widely used DeiT-B and DinoV2 models across various hardware (See Tab.
-3) evaluated on the ImageNet1K [8] dataset.
-These plots demonstrate the relationship between workload size and latency can be variable and non-linear across a variety of hardware.
-Consequently, it is possible to achieve large latency reductions without removing too many tokens in many cases.
-Note that the latency axis is not necessarily indexed from 0 for all plots.
+Other approaches like DynamicViT [36] can yield better accuracy than training-free methods, but require training on server-grade hardware.
+[…]
 We address three shortcomings of existing techniques.
@@
 1 demonstrates the diversity of latency-workload trends across devices and workload sizes.
-As a result, there is room to optimize efficient methods in a hardware aware manner by considering this relationship.
-(2) Some existing efficient methods may require extensive training [42, 30, 5] hinders the deployment of pre-trained models on edge devices.
-(3) Prior work has investigated hardware aware methods for CNN’s [41, 40], the GPU tail effect has not been investigated specifically utilized for transformer models.
-Given these properties of existing token removal methods, this paper focuses on reducing ViT latency by considering hardware characteristics and without requiring training.
+As a result, there is room to optimize efficient methods in a hardware-aware manner by considering this relationship.
+(2) Some existing efficient methods may require extensive training [47, 36, 5], hindering the deployment of pre-trained models on edge devices.
+(3) Prior work has investigated hardware-aware methods for CNN’s [46, 45], but there is little work that shows how to handle or leverage latency-workload behavior for ViT’s.
+Finally, these works lack primary measurements of underlying GPU or kernel behavior.
+Thus, our work focuses on reducing ViT latency by considering ViT latency-workload relationships, and without requiring training.
 This paper has the following contributions:
   1. 1.
-    Introduce the first method for deciding a hardware-aware ViT token pruning schedule for model and device pairs.
+    We identify and profile factors that can affect ViT latency-workload relationships.
   2. 2.
-    Develop a novel training-free token pruning mechanism.
-For similar latency (within 5.2%) across various hardware and workload sizes, we achieve 0.46-43.7 higher accuracy than ToMe, a popular state-of-the-art method.
+    We introduce the first method using latency-workload relationships for deciding ViT token pruning schedules.
+  3. 3.
+    We develop a novel training-free token pruning mechanism.
+For similar latency (within 5.2%) across various hardware and workload sizes, we achieve 0.46 to 43.7 percentage points higher accuracy than ToMe, a state-of-the-art method.
 2 Background and Related Work
-In this section, we review related work on model acceleration and efficient methods for Vision Transformer (ViT) models.
+In this section, we review related work on model acceleration and efficient methods for vision transformers (ViT).
 Some post-processing or fine-tuning methods such as quantization [21] are compatible with token sparsification techniques such as our method.
@@
 State-of-the-art models such as DINOv2-G [30] have over 1 billion parameters.
-It is important to address the efficiency of these models when deploying on edge devices.
-There are many existing model acceleration techniques for ViT’s such as quantization [18], knowledge distillation [39], low-rank factorization [6], and optimized kernels [7] for attention operations.
-These methods often aim to reduce the complexity of the transformer attention mechanism, or learn a compressed representation that has similar effectiveness to a higher-performing model or representation.
-One approach to reduce the latency of transformer models is token sparsification, removing less-important tokens [1, 15, 31, 30] to reduce workload size.
+It is important to address the efficiency of ViT models when deploying on edge devices.
+There are many existing model acceleration techniques for ViT’s such as quantization [21], knowledge distillation [44], low-rank factorization [6], and optimized kernels for attention [8].
+In general, these techniques either remove redundant information or reduce network layer compute.
+One approach related to redundant information removal is token pruning (sparsification) [1, 17, 37, 36].
 These methods are easily applied to a variety of ViT architectures [39, 30, 10].
-A potential benefit of sparsification approaches is that they do not necessarily require any training or fine-tuning [36, 1, 2] .
-Some methods [30, 42, 5, 37] may require significant training time, e.g.
+A potential benefit of pruning approaches is that they do not necessarily require any training or fine-tuning [41, 1, 2].
+However, some methods [36, 47, 5, 42] may require significant training time, e.g.
 100+ epochs, to recover more accuracy and yield latency reductions.
 Training poses a high barrier to application; training-free methods are more accessible [1, 41].
-2.2 The GPU Tail Effect
-The relationship between latency and workload size can be non-linear as seen in Fig.
+2.2 Latency-Workload Relationship
+The thesis of token pruning is that by decreasing workload (tokens), we can reduce latency.
+However, the relationship between latency and workload size can be non-linear as seen in Fig.
 1.
-Large changes in latency are caused by the GPU Tail Effect [40], arising due to workload partitioning and dispatching to the cores of a device’s GPU [23].
-If, at any layer, the GPU cannot process the entire workload at once another wave of computation will be launched.
-A partially full wave costs the same as a full wave.
-The overhead introduced by a partial wave accumulates on a layer-wise basis.
-Consequently, a minor change in workload size may lead to disproportionately large change in latency since the overhead of partial waves is reduced across all layers.
-For example, on the NVIDIA AGX Orin, removing one token can change latency by up to 33% (Fig.
-1).
-Prior work has exploited the tail effect to guide efficient methods for machine learning models [41, 40], particularly in the context of convolutional neural network (CNN) channel pruning [13].
-These methods often reduce latency with negligible accuracy degradation.
-However, ViT token pruning mechanisms are fundamentally different from CNN channel pruning approaches due to architectural differences between CNN’s and ViT’s.
-Therefore, we hypothesize that the GPU tail effect can lower latency for ViT’s, and recognize that new techniques will be needed.
-Our method validates this hypothesis.
-3 Token Pruning with Consideration of the GPU Tail Effect
+This relationship can stem from how workloads are dispatched to the GPU by an ML framework, framework overhead [11], and kernel design decisions [27].
+Tab.
+1 illustrates primary causes of kernel inefficiency.
+# Cause Diagnostic Metrics
+1 Launch Config Occupancy, Grid Size
+2 Memory Usage Memory Bandwidth
+3 Instruction Usage Math Pipe Stalls, Instructions
+Table 1: Primary causes for kernel inefficiency, with associated metrics [27, 16]. Note these causes can co-occur.
+An example of Cause 1 occurs where a kernel grid size is chosen such that a partial wave of computation must be launched on the GPU — this can lead to a phenomenon termed the GPU tail effect [28].
+A partially full wave will incur the same latency cost as a full wave, and this can accumulate on a layer-wise basis.
+Consequently, a minor change in workload size may lead to a disproportionately large change in latency, since the overhead of partial waves accumulates across layers.
+For example, on the NVIDIA AGX Orin [24], removing one token can decrease latency by up to 33% (Fig.
+1.b).
+There are many factors at play that affect latency and kernel behavior.
+Hardware differences, version-to-version ML framework changes, and the use of closed-source backend libraries like cudNN [7] and cuBLAS [26] make it difficult to model or predict neural network latency behavior.
+To illustrate this difficulty, we show more examples in Sec.
+3.1.
+Prior work has exploited the tail effect, which is related to Cause 1, to guide pruning methods for convolutional networks [46, 45].
+Kernel-based optimization methods such as FlashAttention [8] may attempt to choose kernel launch configurations that maximize occupancy.
+Quantization techniques address Cause 2 by using data types that require fewer bits than 32-bit floating point for network parameters.
+Cause 3 can be addressed by choosing low-level operators that might be faster at the cost of precision, or vice-versa [27].
+In this work, we address token pruning in the context of ViT models.
+Previous work frames CNN channel pruning in the context of Cause-1 problems, specifically the GPU tail effect.
+However, ViT token pruning mechanisms are fundamentally different from previous CNN channel pruning [14] approaches due to architectural differences between CNN’s and ViT’s.
+We hypothesize and later demonstrate that latency-workload relationships can also be leveraged to make better token pruning decisions for ViT’s.
+3 Token Pruning with Consideration of Latency and Workload Size
 In Sec.
 2.1 and Sec.
-2.2, we discussed the convenience factor of training-free token pruning methods and how previous work leveraged the GPU tail effect for efficiency benefits.
-Therefore, we choose two design goals for our method: (1) require no training or fine-tuning of a pre-trained ViT model; and (2) achieve more and/or better accuracy-latency tradeoffs by pruning tokens according to the GPU tail effect.
+2.2, we discussed the convenience of training-free token pruning methods and how previous work considered latency and workload size relationships for efficiency benefits.
+Therefore, we set two design goals: (1) require no training or fine-tuning of a pre-trained ViT model; and (2) achieve more and/or better accuracy-latency tradeoffs by pruning tokens according to these relationships.
 Tab.
 2 illustrates qualitative differences between our work and others as a result of these goals.
-Method Training Free? Hardware Aware?
+Method Training Free Hardware Aware
 Ours
 ToMe [1]
-Zero-TPrune [36]
+EViT [19]
 Top-K [13],[19]
 DynamicViT [36]
-STViT [5]
 TPS [42]
-Table 1: Qualitative differences between our pruning approach and similar techniques.
+Table 2: Qualitative differences between our pruning approach and similar techniques. We consider ViT latency-workload relationships in our pruning, making us hardware aware.
 As depicted in Fig.
-2, our token pruning approach consists of two main parts.
+3, our token pruning approach consists of two main parts.
 First, we determine a pruning schedule for a particular model and device pair — this includes identifying the number of tokens to remove and the layers at which to remove them.
@@
 In Sec.
-3.1 we show how to decide the number of tokens to prune based on the GPU tail effect.
+3.1 we show how to decide the number of tokens to prune based on the latency-.
 Next, in Sec.
@@
 3.3 describes our token pruning mechanism which is used at inference time.
-Figure 2: Illustration of our method to decide a pruning schedule (left) and how we prune according to the schedule at inference time (right).
+Figure 2: Latency-workload characteristics of attention operators in PyTorch [34].
+Flash, MemEfficient, and Math are optimized, while Vanilla is not.
+Median latency was measured over 100 runs for each token count.
+All measurements had an IQR $<$ 1us.
+The two annotated latency changes of MemEff are discussed in Sec.
+3.1.
+Figure 3: Illustration of our method to decide a pruning schedule (left), and how we prune according to the schedule at inference time (right).
 3.1 Deciding a Number of Tokens to Prune for our Pruning Schedule
-Workload size definition.
-Before explaining how to leverage the GPU tail effect, we first define workload size.
-We model ViT workload size as being proportional to the number of input tokens $n$, embedding dimension $d$, and batch size $b$ in Equation (1).
-This is because the runtime complexity of the self-attention mechanism is $O(n^{2}\times d)$ [35]; thus the workload size is proportional to $n$ and $d$.
-The batch size $b$ is simply a multiplier for workload size.
-Transformer models are (partially) parameterized by a particular choice of $d$.
-For DeiT-S, $d$ is $384$; for DinoV2-G, $d$ is $1536$.
+ViT workload size.
+Before explaining our method we discuss the core operator of ViT models, the attention mechanism [40].
+ViT models are parameterized (partially) by an embedding dimension $d$, and their inputs have batch size $b$ and a number of tokens $n$.
+The input size, or workload, for attention is a tensor with shape $b\times n\times d$.
 Fig.
 1 illustrates the workload size vs latency relationships by varying $n$, $d$, and $b$ across a set of devices.
-Eq.
-1 describes the relationship in terms of the same quantities:
-$\text{Workload Size}\propto n\times d\times b.$ (1)
-We measure and leverage the GPU tail effect’s characteristics for a particular workload size and device to expand the set of (better) accuracy-latency tradeoffs for token pruning.
-Measuring the GPU tail effect rather than predicting it.
-Prior work has modeled GPU tail effect behavior for a convolutional neural network with some assumptions [17]: (1) The latency staircase pattern increases monotonically; (2) The latency difference between steps is uniform.
-Our experiments indicate that these assumptions may not hold for certain workload sizes using transformer networks.
-First, model latency may actually increase as workload size decreases for certain workload sizes.
-We demonstrate this later in Sec.
-4.
+This wide variety of behavior across devices and workload sizes leads us to consider how ViT latency-workload relationships can be reasonably measured or modeled.
+We now explain underlying reasons for latency-workload non-linearity, and how we decide to measure it.
+Measuring latency behavior effects vs.
+predicting them.
+Inm As previously mentioned, the GPU tail effect is one phenomenon that arises due to suboptimal kernel grid size choice.
+Prior work has modeled [20] or utilized [46, 45] GPU tail effect behavior with respect to convolutional neural networks CNN’s.
+Other work has noted the (sometimes drastic) effect of latency overhead from ML frameworks [11].
+Yu et al.
+[45] notes that the latency of CNN’s increases in discrete jumps, and in all other workload size intervals latency remains the same.
+In our experiments, we find that internal ML framework operator selection can lead to a range of latency behavior.
+This claim for CNN’s does not hold in our evaluations for ViT’s.
 Fig.
-1 (a) demonstrates this phenomenon for DeiT-B on the TX2.
-This anomaly was due to PyTorch [27] launching a different set of kernels that had higher latency for one set of workload sizes.
-Second, the latency difference between steps can be nonuniform.
-The distance between steps varies across workload sizes as seen in Fig.
-1 (a-c).
-Considering these observations, we empirically measure the GPU tail effect due to the difficulty in predicting or modeling its behavior.
-The next part describes the necessary measurements.
-Assigning utility to tokens for deciding a pruning schedule.
-Given a pre-trained model $M$ with a number of input tokens $N$, we frame selection of the number of tokens to prune $R$ as a multi-objective optimization problem between latency gains and accuracy degradation.
-This is the offline computation related to selecting our pruning schedule for $M$ on a target device, as seen in Fig.
-2.
+2 visually illustrates the latency-workload characteristics of various attention operators in PyTorch [32] with $b$=1 and varying $n$ for a DinoV2-G [30] sized attention operation.
+This was measured on an NVIDIA RTX 3090 Ti GPU.
+Additionally, we find that the tail effect is not always the primary factor for drastic latency changes, as was described in previous work.
+We illustrate these findings with measurements of metrics mentioned in Sec.
+2.2 and their co-occurrence with latency behavior.
+In one example from Fig.
+2, the MemEff attention operator features a $\sim$40latency increase from 128 to 130 tokens that is correlated with a $\sim$40increase in pipeline stall or wait time (Cause 3).
+But, the kernel grid size remained the same, indicating that the tail effect is not the primary cause of this latency increase (Cause 1).
+A second example is the $\sim$13latency decrease of MemEff from 132 to 134 tokens is correlated with a 125% increase in kernel grid size (Cause 1) and a 34% decrease in pipeline stall time (Cause 2).
+These effects seem to be downstream of a different kernel being chosen due to internal heuristics.
+Importantly, we note latency can even decrease as workload increases, due to underlying kernel, hardware, and framework behaviors.
+Considering these observations, we empirically measure the latency-workload relationship due to the difficulty in predicting or modeling its behavior.
+The next section describes how we perform this measurement.
+Ranking token importance.
+Given a model $M$ with a number of input tokens $N$, we frame choosing the number of tokens to prune $R$ as a multi-objective optimization problem between latency gains and accuracy degradation.
+This is the offline computation for selecting a pruning schedule for $M$ on a target device, as seen in Fig.
+3.
 First, we measure latency $L(n)$ and accuracy degradation $A(n)$ of $M$ on the target device for each number of tokens to keep $n\in[1,2,...,N]$.
-Measuring latency of $M$ is performed with a grid search across $n$.
-For estimating accuracy degradation, we need a heuristic that does not depend on our pruning schedule selection.
+Measuring latency and accuracy of $M$ is performed with a grid search across $n$.
+We note that in some cases this can be expensive — however we find in many cases it is sufficient to perform the grid search with a stride $>1$ to save time.
+$A(n)$ only needs to be measured once for a particular $M$, and does not need to be done on the target device.
+It is only necessary to record latency $L(n)$ on the target device.
+For accuracy degradation, we need to define $A(n)$ in a way that does not depend on our pruning schedule selection.
 Furthermore it is useful to underestimate the accuracy of $M$ so our selection algorithm is hesitant to remove too many tokens, which can degrade accuracy significantly.
 A simple proxy for the accuracy of each $n$ is to apply random token removal after the first layer on $M$ [31].
-It is assumed that any token pruning method should be better than random token removal since random token pruning does not consider token information content at all.
-Furthermore, pruning at the first layer will degrade accuracy more than pruning later in the network.
+It is assumed that any token pruning method should be better than random token removal since random token pruning does not consider token information content at all [42].
+Furthermore, pruning at the first layer will degrade accuracy more than pruning later in the network [3].
 Therefore random token pruning is a suitable heuristic for estimating accuracy degradation.
+Figure 4: Example offline computation to decide the number of tokens prune.
+We find $n=N_{\text{keep}}$ with maximum utility and by extension the number of tokens to prune $R$ per Eq.
+3.
+Configuration: $\alpha$=0.5, $M$=DinoV2-G, and AGX Orin.
 […]
-$\text{U}_{\text{A}}(n)=\frac{A(n)}{\text{max}\,A(n)}$ (2)
-$\text{U}_{\text{L}}(n)=1-\frac{L(n)}{\text{max}\,L(n)}$ (3)
+$\text{U}_{\text{A}}(n)$ $=\cfrac{A(n)}{\text{max}\,[\,A(n)\,]}$ (1)
+$\text{U}_{\text{L}}(n)$ $=1-\cfrac{L(n)}{\text{max}\,[\,L(n)\,]}$ (2)
 […]
 […]
-Fig.
-3 illustrates an example of the plots of $L(n)$, $A(n)$ and the overall utility $U(n)$.
-With $N=257$, we follow Eq.
-4 to obtain $N_{\text{keep}}=91$ and $R=257-91=166$.
-As part of our evaluation, we measure the effect of different values of $\alpha$ (Sec.
-4.2).
-Figure 3: Example offline computation to decide the number of tokens prune.
-We find $n=N_{\text{keep}}$ with maximum utility and by extension the number of tokens to prune $R$ per Eq.
-4.
-Configuration: $\alpha$=0.5, $M$=DinoV2-G, and AGX Orin.
 3.2 Pruning Schedule: Deciding Layers at which to Prune
-Now we explain the layers to which our pruning mechanism will be applied. This is the final step of the offline computation illustrated in Fig. 2.
-Our schedule prunes all $R$ tokens at one layer, fairly early in the network.
-This differs from other methods such as ToMe [1], Top-K [12, 16], and DynamicViT [30] that progressively prune tokens from layer to layer.
-Our design is based on two observations supported by our evaluation.
-First, on small, edge-grade workloads, repeated pruning operations themselves can introduce non-trivial latency (Sec.
+Now we explain the layers to which our pruning mechanism will be applied.
+This is the final step of the offline computation illustrated in Fig.
+3.
+Our schedule prunes all $R$ tokens at one layer, early in the network.
+This differs from other methods such as ToMe [1], Top-K [13, 19], and DynamicViT [36] that progressively prune tokens.
+Our design is based on two observations supported by our evaluation: First, on small workloads, repeated pruning operations themselves can introduce non-trivial latency (Sec.
 4.5).
-Second, the GPU Tail Effect will lead to improved latency for each layer after which pruning has occurred — earlier pruning means more layers will benefit (Sec.
+Second, latency benefits compound with each layer after pruning occurs — earlier pruning means more layers will benefit (Sec.
 4.3).
-Fig.
-4 illustrates the difference between single-layer and progressive pruning.
-Our approach attempts to place the token pruning at or near the first 25% of layers, similar to the first pruning layer of DynamicViT [30].
+Compared with progressive pruning methods such as ToMe [1] that prune at every layer, our approach attempts to place the token pruning at or near the first 25% of layers, similar to the first pruning layer of DynamicViT [36].
 For DinoV2-G, we prune at the 10th layer, and for DeiT models, we prune at the 4th layer.
@@
 4.3.
-Figure 4: Illustration of the difference between our pruning mechanism and Token Merging [1] ToMe increases latency of DeiT-S for single batch inference on a Jetson TX2 [24] due to the overhead of pruning at every layer.
-Our method is able to yield a latency reduction by pruning fewer tokens at once earlier in the network.
+As an example of the overall process,  Fig.
+4 illustrates an example of the plots of $L(n)$, $A(n)$ and the overall utility $U(n)$.
+With $N=257$, we follow Eq.
+3 to obtain $N_{\text{keep}}=91$ and $R=257-91=166$.
 3.3 Token Pruning Method
@@
 This pruning mechanism is applied during inference (Fig.
-2).
+3).
 The pruning schedule, computed offline, is an input to our token pruning mechanism.
-In terms of design goals, we want our pruning mechanism to require (1) no training or finetuning on $M$, and (2) to introduce minimal overhead.
+In terms of design goals, we want our pruning mechanism to require (1) no finetuning on $M$ and (2) to introduce minimal overhead.
 First, requiring no fine-tuning or training is a convenience for practitioners, and allows pre-trained models to be deployed readily.
@@
 The primary design choice of pruning techniques is choosing a method to rank the importance of tokens.
-Similarly to other pruning techniques [3, 14], we rank the importance of each token $i$ by measuring the quantity of attention received from all other tokens.
-This is analogous to summing along the columns of the attention matrix in a transformer encoder layer.
-Our method incorporates two aspects of importance, one from the $A_{\text{H}}$ matrix and one from the $V$ matrix.
-For attention, we proceed as follows.
-Recall that the core attention mechanism of transformers [35] is parameterized by the $Q$, $K$, and $V$ matrices.
-The attention matrix (including softmax) of a transformer encoder layer with $N$ tokens and $H$ heads is given by $A_{\text{H}}\in\mathbb{R}^{H\times N\times N}$  [35].
-To convert the softmax attention matrix, $A_{\text{H}}$, into tokenwise importance scores we apply a reduction along the along the head dimension, $H$.
-We find the maximum along the head dimension for $A_{\text{H}}$.
-We found that applying max marginally increased accuracy compared to mean or sum.
-This reduction yields a new matrix $A\in\mathbb{R}^{N\times N}$.
-Now, summing across rows of $A$ yields the quantity of attention each token $i$ receives, similar to a component of EViT’s pruning mechanism [16] or Top-K [12].
-This yields an importance score $\text{A}_{\text{S}}[i]\in\mathbb{R}^{N}$ for each token $i$.
-$V_{\text{S}}$ Feature Choice Feature ImageNet1K Accuracy (%) Q 79.74 K 79.77 V 79.81
-$V_{\text{S}}$ Head Reduction Operator Feature ImageNet1K Accuracy (%) mean 79.72 sum 79.70 max 79.81
-Table 2: Token pruning design ablations using for our $V_{S}$ term our schedule on DeiT-S-dist for batch-size 2. In our code we use torch.amax [29] instead of max to handle the batch dimension.
+Similarly to other pruning techniques [3, 15], we rank the importance of each token by measuring the quantity of attention received from all other tokens.
 Figure 5: Illustration of our token pruning method for a single transformer encoder layer at inference time.
@@
 Pruned tokens are aggregated into a single ‘garbage bin’ or inattentive token to retain information content as in EViT [19].
+Our method incorporates two aspects of importance, one from the $A_{\text{H}}$ matrix and one from the $V$ matrix.
+For the $A_{\text{H}}$ term, we proceed as follows.
+Recall that the core attention mechanism of transformers [40] is parameterized by the $Q$, $K$, and $V$ matrices.
+The attention matrix (including softmax) of a transformer encoder layer with $N$ tokens and $H$ heads is given by $A_{\text{H}}\in\mathbb{R}^{H\times N\times N}$  [40].
+To convert the softmax attention matrix $A_{\text{H}}$ into tokenwise scores we: (1) reduce the head dimension $H$ with amax [35] (2) sum along the row $N$ dimension (3) normalize scores by dividing by the max.
+This transforms $A_{\text{H}}$ into tokenwise scores $\text{A}_{\text{S}}\in\mathbb{R}^{N}$, similar to a component of EViT’s pruning mechanism [19] or Top-K [13].
 We then consider a second importance term based on the $V\in\mathbb{R}^{H\times N\times C}$ matrix.
-This term assigns higher scores to tokens whose entry in the $V$ matrix has a large magnitude compared with other tokens.
-Adding this term can marginally improve accuracy.
-Similarly to the attention matrix, we find the maximum along the $H$ dimension [29] — we do a set of ablations summarized in Tab.
-2.
-Post reduction, we have $V^{\prime}\in\mathbb{R}^{N\times C}$ and we have:
-$\text{V}_{\text{S}}[i]$ $=\text{softmax}(\sum^{C}_{c=1}V^{\prime}[i,c])\in\mathbb{R}^{N}.$ (5)
-We add both of these scores to yield our importance score for each token $i$:
-$\text{Importance}[i]$ $=\text{A}_{\text{S}}[i]+\text{V}_{\text{S}}[i].$ (6)
+This term assigns higher scores to tokens who have high weighting in the $V$ matrix, which marginally increased accuracy.
+We perform a similar operation for this term: (1) reduce the $H$ dimension of $V$ with amax [35] (2) sum along the feature dimension $C$ (3) normalize scores with softmax.
+The resulting score vector is $V_{\text{S}}\in\mathbb{R}^{N}$, which we add to $\text{A}_{\text{S}}$ for our overall score:
+Token Importance $=\text{A}_{\text{S}}+\text{V}_{\text{S}}.$ (4)
 Finally, we borrow from EViT [19] and instead of discarding the pruned tokens, we create a new “inattentive” token based on the averaged features of all pruned tokens and append it to the set of kept tokens.
@@
 5 illustrates the token pruning process for a ViT encoder block or layer.
-Note that we do not prune specialized tokens such as the CLS [9] or DIST [33] tokens.
-These tokens serve additional purposes beyond representing the input.
 4 Evaluation
@@
 3) and across various workload sizes.
+We primarily focus on smaller workloads, since these are representative of edge deployment use-cases [18].
 To justify our design decisions, we investigate our $\alpha$ hyperparameter and pruning location choice for our schedule selection and pruning mechanism in Sec.
@@
 Particularly we show that some existing methods actually increase latency rather than reduce it.
+Device GPU CUDA Max Power
+Cores Cores (W)
+TX2 [29] 2 256 15
+Orin [24] 14 1792 40
+A100 [25]  ✝ 108 6192 300
+Table 3: Summary of device hardware information.
+We evaluate our method on two edge devices (TX2 and AGX Orin) and one server-grade system in this work.
+✝ The A100 power consumption is the power consumption only for the GPU, not the entire system.
 4.1 Experimental Setup
@@
 3).
-These devices feature various generations of NVIDIA GPU architectures, and thus feature different levels of computing power.
 We use NVIDIA TX2 [29] and NVIDIA AGX Orin [24] development boards which are edge devices designed for machine learning.
 Additionally, evaluation is done on a server-grade system with an NVIDIA A100 GPU [25] to demonstrate our method can generalize across a wide range of hardware, even though we design our method for edge devices.
-Experiments performed on the TX2 and AGX Orin were performed with fixed CPU and GPU clock rates for reproducibility and consistency.
+Experiments performed on the TX2 and Orin were performed with fixed CPU and GPU clock rates for reproducibility and consistency.
 The A100 system clocks were unable to be locked because the servers machines are university-wide resources.
-Device RAM (GB) GPU Cores CUDA Cores Max Power (W)
-TX2 [24] 8 2 256 15
-AGX Orin [21] 32 14 1792 40
-A100 [22]  ✝ 80 108 6192 300
-Table 3: Summary of device hardware information.
-We evaluate our method on two edge devices (TX2 and AGX Orin) and one server-grade system in this work.
-✝ The A100 power consumption is the power consumption only for the GPU, not the entire system.
 In order to measure latency for evaluation, we use the PyTorch benchmark module [33].
-All accuracy measurements were performed on the A100 system using a batch size of 512.
-Conveniently, accuracy only needs to be measured once for a particular pre-trained model $M$.
-It is only necessary to record the latency $L(n)$ on a target device for a particular workload size in order to capture GPU tail effect behavior.
+All accuracy measurements were performed on the A100 system with batch size 512 on the evaluation split of ImageNet1K.
 In subsequent experiments, we decide the pruning hyperparameters of each token pruning method for fair comparison with our method.
-For reproducibility, our code is open source. We include our method, and modifications to the original ToMe codebase in order to support DinoV2 pre-trained models.
+For reproducibility, our code is open source.
 4.2 Pruning schedule ablation study across $\alpha$ (utility weight of latency/accuracy)
 In 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect.
+Device $\alpha$ R Acc. Median
+Range Loss Latency (ms)
+Orin [0.1, 0.3]0.10.3[0.1,\;0.3][ 0.1 , 0.3 ] 139 1.07 (-27.8%) 112.2
+[0.4, 0.5]0.40.5[0.4,\;0.5][ 0.4 , 0.5 ] 166 2.04 (-33.0%) 104.2
+[0.6, 0.9]0.60.9[0.6,\;0.9][ 0.6 , 0.9 ] 193 4.82 (-35.1%) 100.9
+A100 [0.1, 0.4]0.10.4[0.1,\;0.4][ 0.1 , 0.4 ] 73 0.25 (-4.70%) 28.77
+[0.5, 0.9]0.50.9[0.5,\;0.9][ 0.5 , 0.9 ] 139 1.07 (-6.34%) 28.27
+Table 4: Accuracy vs latency tradeoffs for the computed number of tokens to prune $R$ for various $\alpha$ of our method. Latency percent change is with respect to baseline DinoV2-G inference time.
 The hyperparameter of this algorithm is $\alpha$, which determines the relative weighting of the accuracy and latency utilities.
@@
 Setting to $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens.
-To understand the performance spectrum, we perform an ablation study.
+To understand the performance spectrum, we performed an ablation study.
 The data is in Tab.
-4 and illustrated in Fig.
-6.
-Device $\alpha$ Range Computed $R$ Accuracy Median
-(# tok. to remove) Degradation Latency (ms)
-AGX Orin 0.1-0.3 139 1.07 112.2 (-27.8%)
-0.4-0.5 166 2.04 104.2 (-33.0%)
-0.6-0.9 193 4.82 100.9 (-35.1%)
-A100 0.1-0.4 73 0.25 28.77 (-4.70%)
-0.5-0.9 139 1.07 28.27 (-6.34%)
-Table 4: A companion table for Fig.
-6 that shows the accuracy vs latency tradeoffs for the computed number of tokens to prune $R$ for various $\alpha$.
-Latency percent change is computed with respect to the unmodified DinoV2-G inference time.
-Figure 6: Computed pruning rate $R$ from our offline pruning schedule versus various $\alpha$.
-Note that the selection of $R$ is relatively invariant to changes in $\alpha$.
-Configuration: $M$ = DinoV2-G and batch size = 2.
+4.
 Evaluation on the AGX Orin for $\alpha>0.5$ yielded $R=193$ (over 75% of tokens pruned), which degraded accuracy by $\sim$2.8for 2.1% less latency.
@@
 5 illustrates latency and accuracy tradeoffs for various pruning locations of DinoV2-G.
-Batch Size Pruning Accuracy Median
-Layer Degradation Latency (ms)
-1 1/40 3.19 68.4
+Batch Pruning Acc. Median
+Size Layer Loss Latency (ms)
+1     01/40 3.19 68.4
 10/40 1.07 81.3
@@
 30/40 0.49 104.4
-2 1/40 7.08 79.1
+2     01/40 7.08 79.1
 10/40 2.04 104.7
@@
 That is layer 3 for DeiT-B and DeiT-S, and layer 10 for DinoV2-G.
+Device Batch Model & Acc. Median
+Size Method Loss Latency (ms)
+TX2 2 DeiT-B -(00.0%) 215.0
+w/ Top-K 2.44 (-33.5%) 143.0
+w/ ToMe 1.63 (-33.7%) 142.7
+w/ DyViT ✝ 0.57 (-33.2%) 143.7
+w/ Ours 1.16 (-32.1%) 146.0
+TX2 2 DeiT-S -(00.0%) 68.92
+w/ Top-K 4.30 (-27.0%) 50.32
+w/ ToMe 2.47 (-23.1%) 52.98
+w/ DyViT ✝ 0.46 (-26.2%) 50.88
+w/ Ours 1.24 (-28.3%) 49.44
+Orin 2 DinoV2-G -(00.0%) 155.5
+w/ Top-K 45.66 (-32.9%) 104.4
+w/ ToMe 6.96 (-32.9%) 104.4
+w/ Ours 2.04 (-33.1%) 104.1
+A100 4 DinoV2-G -(00.0%) 40.53
+w/ Top-K 13.31 (-16.7%) 33.76
+w/ ToMe 6.96 (-16.6%) 33.79
+w/ Ours 2.04 (-20.1%) 32.37
+Table 6: Comparison with existing methods for similar latency across device and workload sizes.
+We tune the pruning hyperparameters of the compared methods so they achieve similar latency to our method.
+✝DynamicViT (DyViT) requires training.
+Other evaluated methods are training-free.
 4.4 Accuracy-latency tradeoffs of our method
@@
 Unsurprisingly, DynamicViT is able to retain higher accuracy than all methods, since it requires training.
-Device Batch Model and Accuracy Median
-Size Method Degradation Latency (ms)
-TX2 2 DeiT-B + Top-K 2.44 143.0 (-33.5%)
-DeiT-B + ToMe 1.63 142.7 (-33.7%)
-DeiT-B + Ours 1.16 146.0 (-32.1%)
-DeiT-B + DyViT ✝ 0.57 143.7 (-33.2%)
-DeiT-B 215.0 -(00.0%)
-TX2 2 DeiT-S + Top-K 4.30 50.32 (-27.0%)
-DeiT-S + ToMe 2.47 52.98 (-23.1%)
-DeiT-S + Ours 1.24 49.44 (-28.3%)
-DeiT-S + DyViT ✝ 0.46 50.88 (-26.2%)
-DeiT-S 68.92 -(00.0%)
-AGX Orin 2 DinoV2-G + Top-K 45.66 104.4 (-32.9%)
-DinoV2-G + ToMe 6.96 104.4 (-32.9%)
-DinoV2-G + Ours 2.04 104.1 (-33.1%)
-DinoV2-G 155.5 -(00.0%)
-A100 4 DinoV2-G + Top-K 13.31 33.76 (-16.7%)
-DinoV2-G + ToMe 6.96 33.79 (-16.6%)
-DinoV2-G + Ours 2.04 32.37 (-20.1%)
-DinoV2-G 40.53 -(00.0%)
-Table 6: Comparison with existing methods for similar latency across device and workload sizes.
-We tune the pruning hyperparameters of the compared methods so they achieve similar latency to our method.
-✝DynamicViT (DyViT) requires training.
-Other evaluated methods are training-free.
 Compared with ToMe and Top-K however, it is shown that our method yields consistently lower accuracy degradation.
@@
 6).
-Figure 7: Illustration of accuracy and latency tradeoffs of surveyed methods with $M$ = DinoV2-G.
-Our pruning schedule selection pruning mechanism generate points that expand the pareto front.
-This indicates that considering the GPU tail effect can provide strictly better tradeoffs than existing methods in some cases.
-Note that the number of tokens removed at each layer ($r$) of Top-K and ToMe is evaluated from $r=5$ to $=8$ in increments of 1.
+Figure 6: Illustration of accuracy-latency tradeoffs of surveyed methods with $M$ = DinoV2-G: (a) batch size=2 on AGX Orin (b) batch size=4 on A100 (c) batch size=4 on AGX Orin.
+Our pruning schedule and mechanism generate points that expand the pareto front.
+The number of tokens removed at each layer ($r$) of Top-K and ToMe is evaluated from $r=5$ to $=8$ in increments of 1.
 Second, we note that for larger pruning rates such as $r=7$ and $r=8$, ToMe and Top-K remove nearly all input tokens by the last layer of DinoV2-G.
 These high pruning ratios yield high accuracy degradation of over 40% in the case of A100 batch-size 4 for ToMe, as seen in Fig.
-7.
-Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we simultaneously achieve higher accuracy and lower latency than progressive pruning approaches.
+6.
+Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we achieve higher accuracy and lower latency than progressive pruning approaches.
 This shows aggressive pruning at early layers can be strictly better than progressive pruning methods for certain workload sizes.
-Our token pruning mechanism and schedule offers more accuracy latency tradeoffs across workload sizes.
 4.5 Low workload size observations
@@
 For certain workload sizes, using a baseline model or a hardware aware method such as ours is strictly better than attempting to use a token removal method with high overhead.
-Device Batch Model and Accuracy Median
-Size Method Degradation Latency (ms)
-TX2 1 DeiT-S + Top-K (r=6) 0.73 43.38 (+18.6%)
-DeiT-S + ToMe (r=6) 0.27 50.70 (+30.3%)
-DeiT-S + Ours (R=69) 0.99 32.12 (-9.06%)
-DeiT-S 35.32 (-0.00%)
-Table 7: Given comparable total pruned tokens (69 to 72), for certain workload sizes a token pruning mechanism may increase latency due to the overhead of performing pruning at multiple layers.
+Device Method Acc. Median
+Loss Latency (ms)
+TX2 Baseline (-0.00%) 35.32
+Top-K (r=6) 0.73 (+18.6%) 43.38
+ToMe (r=6) 0.27 (+30.3%) 50.70
+Ours (R=69) 0.99 (-9.06%) 32.12
+Table 7: Experiment on DeiT-S with batch size 1 that illustrates pruning overhead.
+Given comparable total pruned tokens (69 to 72), for certain workload sizes a token pruning mechanism may increase latency due to the overhead of performing pruning at multiple layers.
 Our pruning method is applied at a single layer to avoid this relatively significant overhead.
@@
 3.2, our choice to perform single-layer pruning rather than progressive pruning allows us to yield a latency reduction.
+5 Limitations
+Our approach to token pruning is designed to achieve better tradeoffs when latency-workload is non-linear.
+This occurs primarily for small workload sizes and batch sizes, which are common contexts in edge deployment.
+For large workloads such as DinoV2-G with batch size 256+, our method is less effective because latency-workload relationships becomes linear.
+Last, though we avoid training, our method requires a grid-search.
+In some contexts, this may be expensive (many tokens, large batch size, large model).
+These contexts are outside the focus of this work, but may be relevant for some practitioners.
 6 Conclusion
-In this work, we investigated how to improve the efficiency of pre-trained vision transformers when deployed on edge devices and provided three primary contributions.
-First, we provided insight into how to determine a token pruning schedule for transformer models by leveraging information about the GPU tail effect.
-Second, a token pruning approach was introduced that is training-free and has low overhead compared with other pruning methods.
-We show that our method advances the accuracy-latency pareto front compared with existing token pruning approaches across workload sizes and devices.
-Furthermore we show that for single batch inference, other methods may actually increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%.
-For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
-6 Limitations
-Our approach to token pruning is designed to achieve better tradeoffs at small workload sizes than other methods.
-This is ideal when inputs are streamed to the model, or if only small batch inference is possible.
-For larger workload sizes, such as running DinoV2-G with batch sizes of 32 or higher, our method may not be ideal.
-7 Future Work
-We demonstrate that single layer pruning can be effective even for earlier pruning layers.
-Learning based approaches may be able to take advantage of the high latency reductions of aggressive, early pruning while maintaining accuracy.
-Another promising direction is investigating the benefits of composing token sparsification and other acceleration methods together.
+In this work, we investigated how to improve the efficiency of vision transformers by considering latency-workload relationships.
+We made three main contributions.
+First, we provide insight on how ViT latency is affected by underlying hardware, machine learning framework, and kernel launch factors.
+Second, we show how to determine a token pruning schedule by leveraging information about latency-workload relationships.
+Third, we introduced a token pruning mechanism that is training-free and low overhead.
+Our method advances the accuracy-latency Pareto front compared with existing token pruning approaches across workload sizes and devices.
+For single batch inference, other methods may increase latency by 18.6-30.3% with respect to baseline, while we reduce it by 9%.
+For similar latency (within 5.2%) across devices, the state-of-the-art, ToMe, achieves 45.8%-85.4% ImageNet1K accuracy, while we achieve 78.6%-84.5%.
+Improving kernel designs to reduce non-linear latency behavior is a promising future direction.

### v3

@@
 Abstract
-This paper investigates how to efficiently deploy vision transformers on edge devices.
+This paper investigates how to efficiently deploy vision transformers on edge devices for small workloads.
 Recent methods reduce the latency of transformer neural networks by removing or merging tokens, with small accuracy degradation.
-However, these methods are not designed with edge device deployment in mind: jthey do not leverage information about the latency vs.
-workload trends to improve efficiency.
-First, we show the latency-workload size relationship is nonlinear for certain workload sizes.
-We consider this relationship to create a token pruning schedule.
-Second, we demonstrate a training-free, token pruning method utilizing this schedule.
-We show that for single batch inference, other methods increase latency by 18.6-30.3% with respect to baseline, while we can reduce it by 9%.
-For similar latency (within 5.2%) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
+However, these methods are not designed with edge device deployment in mind: they do not leverage information about the latency-workload trends to improve efficiency.
+We address this shortcoming in our work.
+First, we identify factors that affect ViT latency-workload relationships.
+Second, we determine token pruning schedule by leveraging non-linear latency-workload relationships.
+Third, we demonstrate a training-free, token pruning method utilizing this schedule.
+We show other methods may increase latency by 2-30%, while we reduce latency by 9-26%.
+For similar latency (within 5.2% or 7ms) across devices we achieve 78.6%-84.5% ImageNet1K accuracy, while the state-of-the-art, Token Merging, achieves 45.8%-85.4%.
 1 Introduction
@@
 Edge devices have become sufficiently powerful, and model miniaturization techniques sufficiently capable, that machine learning (ML) models can be deployed to the network edge for various tasks, including computer vision applications [11, 36].
-However, state of the art performance on various computer vision tasks are often claimed by large vision transformer (ViT) based neural network [10] architectures.
+However, state of the art performance on various computer vision tasks is often claimed by large vision transformer (ViT) based neural network [9] architectures.
 ViT models such as DeiT [37] and DINOv2 [29] are not designed nor miniaturized for edge deployment.
-Additionally, latency is often of critical importance on edge devices, and only small inputs or workloads can be processed [18].
-Prior work has shown that ViT’s have high redundancy that can be exploited for latency reduction benefits [23].
-One general approach is to identify and remove parts of the input (tokens) with low information content.
-This is called token sparsification.
+Additionally, latency (the time required to do a forward pass given a batch of inputs) is often of critical importance on edge devices, and only small inputs or workloads can be processed [17].
+Prior work has shown that ViTs have high redundancy that can be exploited for latency reduction benefits [22].
+One approach involves identifying and removing low-information tokens; this is called token sparsification.
 Training-free token sparsification methods such as Token Merging (ToMe) [1] have been effective at reducing latency of pre-trained models.
@@
 […]
-We address three shortcomings of existing techniques.
+We address three key shortcomings in existing techniques.
 (1) Many existing efficient methods do not consider fine-grained hardware or latency characteristics [1, 18, 40].
@@
 1 demonstrates the diversity of latency-workload trends across devices and workload sizes.
-As a result, there is room to optimize efficient methods in a hardware-aware manner by considering this relationship.
+As a result, there is room to improve efficient methods in a hardware-aware manner by considering this relationship.
 (2) Some existing efficient methods may require extensive training [46, 34, 4], hindering the deployment of pre-trained models on edge devices.
-(3) Prior work has investigated hardware-aware methods for CNN’s [46, 45], but there is little work that shows how to handle or leverage latency-workload behavior for ViT’s.
-Finally, these works lack primary measurements of underlying GPU or kernel behavior.
+(3) Prior work has investigated hardware-aware methods for CNNs [45, 44], but there is little work that shows how to handle or leverage latency-workload behavior for ViTs.
+Finally, these works lack direct measurements of underlying GPU or kernel behavior.
 Thus, our work focuses on reducing ViT latency by considering ViT latency-workload relationships, and without requiring training.
@@
   2. 2.
-    We introduce the first method using latency-workload relationships for deciding ViT token pruning schedules.
+    We propose the first method using latency-workload relationships for deciding ViT token pruning schedules.
   3. 3.
-    We develop a novel training-free token pruning mechanism.
-For similar latency (within 5.2%) across various hardware and workload sizes, we achieve 0.46 to 43.7 percentage points higher accuracy than ToMe, a state-of-the-art method.
+    We design a novel training-free token pruning mechanism.
+For similar latency across various hardware and workload sizes, we achieve 0.46 to 43.7 percentage points higher accuracy than ToMe, a state-of-the-art method.
 2 Background and Related Work
@@
 It is important to address the efficiency of ViT models when deploying on edge devices.
-There are many existing model acceleration techniques for ViT’s such as quantization [21], knowledge distillation [44], low-rank factorization [6], and optimized kernels for attention [8].
+Numerous techniques exist for accelerating ViT models, including: quantization [20], knowledge distillation [43], low-rank factorization [5], and optimized kernels for attention [7].
 In general, these techniques either remove redundant information or reduce network layer compute.
-One approach related to redundant information removal is token pruning (sparsification) [1, 17, 37, 36].
+One approach related to redundant information removal is token sparsification [1, 16, 35, 34].
 These methods are easily applied to a variety of ViT architectures [37, 29, 9].
-A potential benefit of pruning approaches is that they do not necessarily require any training or fine-tuning [41, 1, 2].
+One advantage of sparsification methods is that they often do not necessarily require training or fine-tuning [39, 1].
 However, some methods [34, 46, 4, 40] may require significant training time, e.g.
@@
 2.2 Latency-Workload Relationship
-The thesis of token pruning is that by decreasing workload (tokens), we can reduce latency.
-However, the relationship between latency and workload size can be non-linear as seen in Fig.
+The premise behind token pruning is that reducing the workload (tokens) can decrease latency.
+However, this relationship can be non-linear, as demonstrated in Fig.
 1.
@@
 An example of Cause 1 occurs where a kernel grid size is chosen such that a partial wave of computation must be launched on the GPU — this can lead to a phenomenon termed the GPU tail effect [27].
-A partially full wave will incur the same latency cost as a full wave, and this can accumulate on a layer-wise basis.
-Consequently, a minor change in workload size may lead to a disproportionately large change in latency, since the overhead of partial waves accumulates across layers.
-For example, on the NVIDIA AGX Orin [24], removing one token can decrease latency by up to 33% (Fig.
+A partially filled wave incurs the same latency cost as a fully filled one, and this effect can compound across layers.
+Thus, even minor adjustments in workload size can result in significant latency changes due to the cumulative overhead of partial waves across layers.
+For example, on the NVIDIA AGX Orin [23], removing one token (97 to 96) can decrease latency by up to 33% (Fig.
 1.b).
-There are many factors at play that affect latency and kernel behavior.
-Hardware differences, version-to-version ML framework changes, and the use of closed-source backend libraries like cudNN [7] and cuBLAS [26] make it difficult to model or predict neural network latency behavior.
+There are many factors that affect latency.
+Differences in hardware, changes across ML framework versions, and reliance on proprietary backend libraries like cudNN [6] and cuBLAS [25] complicate the modeling and prediction of neural network latency.
 To illustrate this difficulty, we show more examples in Sec.
@@
 Kernel-based optimization methods such as FlashAttention [7] may attempt to choose kernel launch configurations that maximize occupancy.
-Quantization techniques address Cause 2 by using data types that require fewer bits than 32-bit floating point for network parameters.
+Quantization addresses Cause 2 by employing data types with fewer bits than 32-bit floating point for network parameters.
 Cause 3 can be addressed by choosing low-level operators that might be faster at the cost of precision, or vice-versa [26].
@@
 Previous work frames CNN channel pruning in the context of Cause-1 problems, specifically the GPU tail effect.
-However, ViT token pruning mechanisms are fundamentally different from previous CNN channel pruning [14] approaches due to architectural differences between CNN’s and ViT’s.
-We hypothesize and later demonstrate that latency-workload relationships can also be leveraged to make better token pruning decisions for ViT’s.
+However, ViT token pruning mechanisms are fundamentally different from previous CNN channel pruning [13] approaches due to architectural differences between CNNs and ViTs.
+We hypothesize and later demonstrate that latency-workload relationships can also be leveraged to make better token pruning decisions for ViTs.
 3 Token Pruning with Consideration of Latency and Workload Size
@@
 2.1 and Sec.
-2.2, we discussed the convenience of training-free token pruning methods and how previous work considered latency and workload size relationships for efficiency benefits.
-Therefore, we set two design goals: (1) require no training or fine-tuning of a pre-trained ViT model; and (2) achieve more and/or better accuracy-latency tradeoffs by pruning tokens according to these relationships.
+2.2, we discussed the advantages of training-free token pruning methods and how previous work considered latency and workload size relationships for efficiency benefits.
+Therefore, we set two design goals: (1) require no training or fine-tuning of a pre-trained ViT model; and (2) achieve better accuracy-latency tradeoffs by pruning tokens according to these relationships.
 Tab.
@@
 TPS [40]
-Table 2: Qualitative differences between our pruning approach and similar techniques. We consider ViT latency-workload relationships in our pruning, making us hardware aware.
+Table 2: Qualitative differences between our pruning approach and similar techniques. We consider ViT latency-workload relationships to guide our token pruning.
 As depicted in Fig.
 3, our token pruning approach consists of two main parts.
-First, we determine a pruning schedule for a particular model and device pair — this includes identifying the number of tokens to remove and the layers at which to remove them.
-Second, we develop a training-free pruning technique to remove non-informative tokens during inference time.
+First, we establish a pruning schedule for specific model-device pairs, which involves determining the number of tokens to prune and the layers where pruning occurs.
+Second, we devise a training-free technique for pruning non-informative tokens at inference time.
 In Sec.
-3.1 we show how to decide the number of tokens to prune based on the latency-.
+3.1 we show how to decide the number of tokens to prune based on the latency-workload relationship.
 Next, in Sec.
 3.2 we explain our choice of which layers to prune, completing our offline pruning schedule selection.
-Finally, Sec.
+Sec.
 3.3 describes our token pruning mechanism which is used at inference time.
-Figure 2: Latency-workload characteristics of attention operators in PyTorch [34].
-Flash, MemEfficient, and Math are optimized, while Vanilla is not.
-Median latency was measured over 100 runs for each token count.
-All measurements had an IQR $<$ 1us.
-The two annotated latency changes of MemEff are discussed in Sec.
-3.1.
-Figure 3: Illustration of our method to decide a pruning schedule (left), and how we prune according to the schedule at inference time (right).
+Last, Sec.
+3.4 clarifies qualitative differences between our method and existing approaches.
+[…]
+Figure 3: Illustration of our method to decide a pruning schedule (left) and how we prune according to the schedule at inference time (right), which are discussed in the sectiions shown at the top of the illustration.
 3.1 Deciding a Number of Tokens to Prune for our Pruning Schedule
@@
 Before explaining our method we discuss the core operator of ViT models, the attention mechanism [38].
-ViT models are parameterized (partially) by an embedding dimension $d$, and their inputs have batch size $b$ and a number of tokens $n$.
+ViT models are partially parameterized by an embedding dimension $d$, with inputs characterized by a batch size $b$ and a number of tokens $n$.
 The input size, or workload, for attention is a tensor with shape $b\times n\times d$.
 Fig.
-1 illustrates the workload size vs latency relationships by varying $n$, $d$, and $b$ across a set of devices.
+1 demonstrates how varying $b$, $n$, and $d$ affects the relationship between workload size and latency across various devices.
 This wide variety of behavior across devices and workload sizes leads us to consider how ViT latency-workload relationships can be reasonably measured or modeled.
@@
 predicting them.
-Inm As previously mentioned, the GPU tail effect is one phenomenon that arises due to suboptimal kernel grid size choice.
-Prior work has modeled [20] or utilized [46, 45] GPU tail effect behavior with respect to convolutional neural networks CNN’s.
+As previously mentioned, the GPU tail effect is one phenomenon that arises due to suboptimal kernel grid size choice.
+Prior work has modeled [19] or utilized [45, 44] GPU tail effect behavior with respect to convolutional neural networks CNNs.
 Other work has noted the (sometimes drastic) effect of latency overhead from ML frameworks [10].
 Yu et al.
-[45] notes that the latency of CNN’s increases in discrete jumps, and in all other workload size intervals latency remains the same.
+[44] claims that the latency of CNNs increases in discrete jumps, and in all other workload size intervals latency remains the same.
 In our experiments, we find that internal ML framework operator selection can lead to a range of latency behavior.
-This claim for CNN’s does not hold in our evaluations for ViT’s.
+This claim for CNNs does not hold in our evaluations for ViTs.
 Fig.
-2 visually illustrates the latency-workload characteristics of various attention operators in PyTorch [32] with $b$=1 and varying $n$ for a DinoV2-G [30] sized attention operation.
+2 depicts the latency-workload characteristics of various attention operators in PyTorch [31] with $b$=1 and varying $n$ for a DinoV2-G [29] attention operation.
 This was measured on an NVIDIA RTX 3090 Ti GPU.
@@
 2, the MemEff attention operator features a $\sim$40latency increase from 128 to 130 tokens that is correlated with a $\sim$40increase in pipeline stall or wait time (Cause 3).
-But, the kernel grid size remained the same, indicating that the tail effect is not the primary cause of this latency increase (Cause 1).
+But, the kernel grid size remained the same, indicating that the tail effect (Cause 1) was not the primary cause of this latency increase.
 A second example is the $\sim$13latency decrease of MemEff from 132 to 134 tokens is correlated with a 125% increase in kernel grid size (Cause 1) and a 34% decrease in pipeline stall time (Cause 2).
@@
 Ranking token importance.
-Given a model $M$ with a number of input tokens $N$, we frame choosing the number of tokens to prune $R$ as a multi-objective optimization problem between latency gains and accuracy degradation.
+Given a model $M$ with $N$ input tokens, we define the selection of $R$ tokens to prune as a multi-objective optimization problem, balancing latency gains against accuracy degradation.
 This is the offline computation for selecting a pruning schedule for $M$ on a target device, as seen in Fig.
 3.
-First, we measure latency $L(n)$ and accuracy degradation $A(n)$ of $M$ on the target device for each number of tokens to keep $n\in[1,2,...,N]$.
-Measuring latency and accuracy of $M$ is performed with a grid search across $n$.
-We note that in some cases this can be expensive — however we find in many cases it is sufficient to perform the grid search with a stride $>1$ to save time.
-$A(n)$ only needs to be measured once for a particular $M$, and does not need to be done on the target device.
-It is only necessary to record latency $L(n)$ on the target device.
-For accuracy degradation, we need to define $A(n)$ in a way that does not depend on our pruning schedule selection.
+First, we measure latency $L(n)$ of $M$ on the target device for each number of tokens to keep $n\in[1,2,...,N]$.
+Measuring latency and accuracy of $M$ is performed with a grid search across $n$this is demonstrated in Algorithm 1.
+Running times for all configurations we evaluate over are listed later in Sec.
+4.2.
+{minted}
+[fontsize=]python def measure_latency(model, b, N): """ b refers to batch size, N is the number of tokens ’model’ expects """ #L(n), as a dictionary L = for n in range(1, N): # Latency is independent of random inputs # model.d is the embedding size (Sec.
+3.1) x = torch.rand(b,n,model.d) # Benchmark for fixed time latency = bench(model, x) L[n] = latency return L
+Algorithm 1 Offline Workload Latency Measurement
+{minted}
+[fontsize=]python def measure_accuracy(model, dataset, N): """ dataset is the eval split, N is the number of tokens ’model’ expects """ #A(n), as a dictionary A = for n in range(1, N): # Running accuracy for n n_acc = 0.0 for image,label in dataset: # Shape (b, N, model.d) x = model.embed(image) # Shape (b, n, model.d) x = random_prune(x, n) y_pred = model.predict(x) n_acc += sum(y_targ == label) A[n] = n_acc / len(dataset) return A
+Algorithm 2 Offline Accuracy Degradation Estimate
+Second, we estimate the accuracy $A(n)$ of $M$ after pruning tokens.
+We need to measure $A(n)$ in a way that does not depend on our pruning schedule selection, however.
 Furthermore it is useful to underestimate the accuracy of $M$ so our selection algorithm is hesitant to remove too many tokens, which can degrade accuracy significantly.
-A simple proxy for the accuracy of each $n$ is to apply random token removal after the first layer on $M$ [31].
+A simple proxy for the accuracy of each $n$ is to apply random token removal after the first layer on $M$ [30], which we refer to as random_prune.
+Algorithm 2 depicts how to compute $A(n)$.
 It is assumed that any token pruning method should be better than random token removal since random token pruning does not consider token information content at all [40].
 Furthermore, pruning at the first layer will degrade accuracy more than pruning later in the network [2].
-Therefore random token pruning is a suitable heuristic for estimating accuracy degradation.
-Figure 4: Example offline computation to decide the number of tokens prune.
-We find $n=N_{\text{keep}}$ with maximum utility and by extension the number of tokens to prune $R$ per Eq.
-3.
-Configuration: $\alpha$=0.5, $M$=DinoV2-G, and AGX Orin.
+Therefore random token pruning is a suitable heuristic for estimating accuracy.
 […]
+The following definitions meet these criteria:
+$\text{U}_{\text{L}}(n)$ $=1-\cfrac{L(n)}{\text{max}\,[\,L(n)\,]}$ (1)
 $\text{U}_{\text{A}}(n)$ $=\cfrac{A(n)}{\text{max}\,[\,A(n)\,]}$ (2)
-$\text{U}_{\text{L}}(n)$ $=1-\cfrac{L(n)}{\text{max}\,[\,L(n)\,]}$ (2)
 […]
@@
 3.2 Pruning Schedule: Deciding Layers at which to Prune
-Now we explain the layers to which our pruning mechanism will be applied.
-This is the final step of the offline computation illustrated in Fig.
-3.
-Our schedule prunes all $R$ tokens at one layer, early in the network.
+We explain where in the ViT our pruning mechanism is applied.
+Our schedule prunes all $R$ tokens at one layer, early in the model.
 This differs from other methods such as ToMe [1], Top-K [12, 18], and DynamicViT [34] that progressively prune tokens.
-Our design is based on two observations supported by our evaluation: First, on small workloads, repeated pruning operations themselves can introduce non-trivial latency (Sec.
-4.5).
-Second, latency benefits compound with each layer after pruning occurs — earlier pruning means more layers will benefit (Sec.
+This choice is based on two observations supported by our evaluation: First, on small workloads, the repeated application of pruning operations can introduce significant latency (Sec.
 4.3).
-Compared with progressive pruning methods such as ToMe [1] that prune at every layer, our approach attempts to place the token pruning at or near the first 25% of layers, similar to the first pruning layer of DynamicViT [36].
-For DinoV2-G, we prune at the 10th layer, and for DeiT models, we prune at the 4th layer.
-This allows us to realize latency reductions according to the GPU tail effect for the remaining 30 layers in DinoV2-G and the 8 remaining layers for DeiT-S and DeiT-B.
-We discuss the effect of pruning location in Sec.
-4.3.
-As an example of the overall process,  Fig.
-4 illustrates an example of the plots of $L(n)$, $A(n)$ and the overall utility $U(n)$.
-With $N=257$, we follow Eq.
-3 to obtain $N_{\text{keep}}=91$ and $R=257-91=166$.
+Second, latency reductions accumulate with each subsequent layer after pruning; thus, pruning earlier allows more layers to benefit from low latency.
+{minted}
+[fontsize=]python def vit_forward(model, x, N, R, L): """ x is an image-like input.
+N is the number of tokens after embedding x.
+R is the number of tokens to prune, and L is the layer at which to prune (Sec.
+3.2).
+""" # Tokenize input image, x has shape (b, N, d) x = model.embed(x) for idx, layer in enumerate(model.layers): # Standard self-attention x, attn, V = layer(x) if idx == L: scores = rank_tokens(attn, V) x = prune_tokens(x, scores, N, R) return model.head(x)
+def rank_tokens(attn, V): """ attn and V are from self-attention.
+attn has shape (b, h, N, N), where h is the number of attention heads.
+V has shape (b, h, N, d / h).
+’keepdim=True’ means the tensor dimension reduced over is not removed, but is kept with length=1.
+""" am = max(attn,dim=1).sum(dim=1,keepdim=True) am /= max(am.transpose(-2,-1)) # V metric (ours) vm = max(V, dim=1).sum(dim=-1,keepdim=True) vm = softmax(vm, dim=1) # Tokenwise scores with shape (b, N, 1) return am + vm
+def prune_tokens(x, scores, N, R): """ x is a tensor of tokens.
+We prune such that N-R tokens remain.
+""" # Sort by score, descending sorted_scores_idx = argsort(scores,dim=1) # Shape (b, N-R-1, d) kept = gather(x, sorted_scores[:,:N-R-1]) # Shape (b, 1, d) inattentive = gather(x, sorted_scores[:,N-R-1:]).mean(dim=1) # Shape (b, N-R, d) return torch.cat([kept, inattentive], dim=1)
+Algorithm 3 Inference-Time Pruning Mechanism
+We perform pruning after the first 25% of ViT layers, akin to the first pruning layer of DynamicViT [34] — this yields latency reduction for the remaining 75% of layers.
+We refer to the index of this pruning layer as $L$.
+Different pruning locations are evaluated in the supplemental work.
 3.3 Token Pruning Method
@@
 3.1 and 3.2 identifies the location and number of tokens to prune.
-Now we give a training-free token pruning mechanism to decide which tokens to prune.
-This pruning mechanism is applied during inference (Fig.
+Now, we give a training-free token pruning mechanism to decide which tokens to prune at inference as in Fig.
+3.
+The offline pruning schedule, consisting of a number of tokens to prune $R$ and the layer index $L$ at which to prune, is an input to our token pruning mechanism.
+The primary design goal for our pruning mechanism is to require no finetuning on $M$ and be lightweight - thus we restrict ourselves to using intermediate computations from the attention operation.
+First, we choose a method to rank the importance of tokens.
+Following prior work [2, 14], we rank token importance by measuring the attention each token receives from all others, utilizing the softmax attention matrix.
+We also incorporate an importance term derived from the $V$ matrix, which marginally increased accuracy.
+Second, we borrow from EViT [18] and instead of discarding pruned tokens, we create a new “inattentive” token based on the features of all pruned tokens, then and append it to the set of kept tokens.
+Information is thus preserved from the pruned tokens, increasing accuracy while reducing the total token count.
+Algorithm 3 depicts a forward pass using our pruning mechanism.
+Inputs from our offline computation, $R$ and $L$, are utilized to rank tokens and decide at which layer to prune.
+rank_tokens ranks each token based on our V matrix importance vm and a standard attention matrix importance term am.
+prune_tokens prunes $R$ tokens using a standard token removal implementation [34].
+3.4 Qualitative Comparison with Pruning and Merging
+Here, we justify why we classify our method as token pruning rather than merging.
+Our method uses the “inattentive“ token from EViT, which the EViT authors consider a hybrid method.
+However, we take the position that the core mechanism of deciding which tokens to prune is an important differentiating factor.
+A majority of pruning-based approaches treat the selection of tokens to prune as a ranking problem, rather than a matching problem as merging methods do.
+Our importance score computation is most similar to ranking-based approaches.
+Thus, we see ourselves primarily as a pruning method, though we could be considered a hybrid between pruning and merging.
+4 Evaluation
+After describing experimental setup (Sec.
+4.1), we characterize our technique via ablation over $\alpha$ and by measuring offline computation costs (Sec.
+4.2).
+Then we compare to the state-of-art ToMe method and relevant baselines (Sec.
+4.3).
+4.1 Experimental Setup
+Hardware: We use three devices with varying characteristics (Tab.
 3).
-The pruning schedule, computed offline, is an input to our token pruning mechanism.
-In terms of design goals, we want our pruning mechanism to require (1) no finetuning on $M$ and (2) to introduce minimal overhead.
-First, requiring no fine-tuning or training is a convenience for practitioners, and allows pre-trained models to be deployed readily.
-Second, keeping overhead to a minimum for small workload sizes is necessary to yield latency reductions as previously mentioned in 3.2.
-The primary design choice of pruning techniques is choosing a method to rank the importance of tokens.
-Similarly to other pruning techniques [3, 15], we rank the importance of each token by measuring the quantity of attention received from all other tokens.
-Figure 5: Illustration of our token pruning method for a single transformer encoder layer at inference time.
-We compute an attention score and V score for each token to estimate importance, then prune the least informative tokens.
-Pruned tokens are aggregated into a single ‘garbage bin’ or inattentive token to retain information content as in EViT [19].
-Our method incorporates two aspects of importance, one from the $A_{\text{H}}$ matrix and one from the $V$ matrix.
-For the $A_{\text{H}}$ term, we proceed as follows.
-Recall that the core attention mechanism of transformers [40] is parameterized by the $Q$, $K$, and $V$ matrices.
-The attention matrix (including softmax) of a transformer encoder layer with $N$ tokens and $H$ heads is given by $A_{\text{H}}\in\mathbb{R}^{H\times N\times N}$  [40].
-To convert the softmax attention matrix $A_{\text{H}}$ into tokenwise scores we: (1) reduce the head dimension $H$ with amax [35] (2) sum along the row $N$ dimension (3) normalize scores by dividing by the max.
-This transforms $A_{\text{H}}$ into tokenwise scores $\text{A}_{\text{S}}\in\mathbb{R}^{N}$, similar to a component of EViT’s pruning mechanism [19] or Top-K [13].
-We then consider a second importance term based on the $V\in\mathbb{R}^{H\times N\times C}$ matrix.
-This term assigns higher scores to tokens who have high weighting in the $V$ matrix, which marginally increased accuracy.
-We perform a similar operation for this term: (1) reduce the $H$ dimension of $V$ with amax [35] (2) sum along the feature dimension $C$ (3) normalize scores with softmax.
-The resulting score vector is $V_{\text{S}}\in\mathbb{R}^{N}$, which we add to $\text{A}_{\text{S}}$ for our overall score:
-Token Importance $=\text{A}_{\text{S}}+\text{V}_{\text{S}}.$ (4)
-Finally, we borrow from EViT [19] and instead of discarding the pruned tokens, we create a new “inattentive” token based on the averaged features of all pruned tokens and append it to the set of kept tokens.
-This allows information from pruned tokens to be preserved in this inattentive token, which EViT found improved accuracy.
-Fig.
-5 illustrates the token pruning process for a ViT encoder block or layer.
-4 Evaluation
-In this section, we demonstrate the effectiveness of our token pruning schedule and pruning mechanism across devices (listed in Tab.
-3) and across various workload sizes.
-We primarily focus on smaller workloads, since these are representative of edge deployment use-cases [18].
-To justify our design decisions, we investigate our $\alpha$ hyperparameter and pruning location choice for our schedule selection and pruning mechanism in Sec.
-4.2 and Sec.
-4.3, respectively.
-Then, we compare with a variety of existing token removal methods such as DynamicViT [36], ToMe [1].
-We use Top-K [19, 13] as a baseline on ImageNet1K [9].
-Sec.
-4.4 demonstrates how our method advances the accuracy vs.
-latency pareto front across devices, workload sizes, and models.
-Finally, we identify some pitfalls of existing token pruning techniques for small workload sizes in Sec.
-4.5.
-Particularly we show that some existing methods actually increase latency rather than reduce it.
+Our technique targets edge workloads, so we use two edge-caliber development boards designed for machine learning: NVIDIA TX2 [28] and NVIDIA AGX Orin [23].
+To assess generalization beyond edge devices, we also use a server-grade NVIDIA A100 GPU [24].
+On the TX2 and Orin we used fixed CPU and GPU clock rates for consistency.
+The A100 system clocks could not be locked because those servers are shared resources.
 Device GPU CUDA Max Power
@@
 ✝ The A100 power consumption is the power consumption only for the GPU, not the entire system.
-4.1 Experimental Setup
-Evaluation is done using three devices with varying hardware characteristics (Tab.
-3).
-We use NVIDIA TX2 [29] and NVIDIA AGX Orin [24] development boards which are edge devices designed for machine learning.
-Additionally, evaluation is done on a server-grade system with an NVIDIA A100 GPU [25] to demonstrate our method can generalize across a wide range of hardware, even though we design our method for edge devices.
-Experiments performed on the TX2 and Orin were performed with fixed CPU and GPU clock rates for reproducibility and consistency.
-The A100 system clocks were unable to be locked because the servers machines are university-wide resources.
-In order to measure latency for evaluation, we use the PyTorch benchmark module [33].
-All accuracy measurements were performed on the A100 system with batch size 512 on the evaluation split of ImageNet1K.
+Model Params (M) Depth
+DeiT-S [37] 21 12
+DeiT-B [37] 86 12
+ViT-L [9, 42] 300 24
+DinoV2-G [29] 1100 40
+Table 4: Characteristics of models we evaluate over. We include DinoV2 as a representative state-of-the-art ViT, while ViT and DeiT are commonly used baselines in prior token pruning work.
+Models: Tab.
+4 summarizes the models used.
+We evaluate common vision transformer models across a variety of scales (21M to 1.1B parameters).
+For DeiT and ViT models we use the TIMM pretrained weights, while we use the DinoV2-G weights from the DinoV2 Github.
+Measurements: In order to measure latency for evaluation, we use the PyTorch benchmark module [32].
+Latency is measured over 16 seconds.
+To be clear, we define latency as the compute time required for a forward pass of a model given a batch of input images.
+Accuracy was measured using the A100 system, with a batch size of 512 on the evaluation subset of ImageNet1K.
 In subsequent experiments, we decide the pruning hyperparameters of each token pruning method for fair comparison with our method.
 For reproducibility, our code is open source.
-4.2 Pruning schedule ablation study across $\alpha$ (utility weight of latency/accuracy)
-In 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect.
-Device $\alpha$ R Acc. Median
+4.2 Characterizations of our technique
+Here we evaluate two of the three design decisions of our method.
+First, we ablate over the hyperparameter $\alpha$ that is used to weight the utility function.
+Second, we measure the cost (time) for offline computation.
+The third decision, selecting the layer $L$ for pruning, is evaluated in supplemental material.
+Pruning schedule ablation study across $\alpha$. In Sec. 3.1 we introduce an algorithm to decide a number of tokens to prune $R$ according to the GPU tail effect.
+[…]
 Range Loss Latency (ms)
@@
 Table 5: Accuracy vs latency tradeoffs for the computed number of tokens to prune $R$ for various $\alpha$ of our method. Latency percent change is with respect to baseline DinoV2-G inference time.
-The hyperparameter of this algorithm is $\alpha$, which determines the relative weighting of the accuracy and latency utilities.
-Setting to $\alpha=1$ prioritizes accuracy, and only a few tokens might be pruned.
-Setting to $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens.
-To understand the performance spectrum, we performed an ablation study.
-The data is in Tab.
-4.
-Evaluation on the AGX Orin for $\alpha>0.5$ yielded $R=193$ (over 75% of tokens pruned), which degraded accuracy by $\sim$2.8for 2.1% less latency.
+The $\alpha$ hyperparameter governs the relative weighting of accuracy and latency utilities in our algorithm.
+Setting $\alpha=1$ prioritizes accuracy, and only a few tokens might be pruned.
+Setting $\alpha=0$ prioritizes latency, and would prune all or nearly-all tokens.
+To decide the value of $\alpha$, we performed an ablation study in  Tab.
+5.
+When evaluated on the AGX Orin with $\alpha>0.5$, we computed $R=193$ (over 75% of tokens pruned), resulting in a $\sim$2.8accuracy drop for a 2.1% latency reduction.
 We consider this to be an unfavorable tradeoff.
 This ablation, in addition to results in Sec.
-4.4, suggest that $\alpha\leq 0.5$ is a good choice.
+4.3, suggest that $\alpha\leq 0.5$ is a good choice.
 Thus we use $\alpha=0.5$ for all evaluations that appear in this work.
@@
 3 when selecting $R$.
-4.3 Pruning location ablation study
-In 3.2 we decide at which layer our pruning mechanism should be applied.
-To provide insight into the potential pruning locations, we performed an ablation study.
+Offline computation time.
+In  Sec.
+3.1, we describe our offline computations to decide a number of tokens to prune in which we utilize a grid search to measure latency and estimate accuracy degradation.
 Tab.
-5 illustrates latency and accuracy tradeoffs for various pruning locations of DinoV2-G.
-Batch Pruning Acc. Median
-Size Layer Loss Latency (ms)
-1     01/40 3.19 68.4
-10/40 1.07 81.3
-20/40 0.58 93.4
-30/40 0.49 104.4
-2     01/40 7.08 79.1
-10/40 2.04 104.7
-20/40 0.97 133.0
-30/40 0.83 160.8
-Table 5: Latency/accuracy tradeoffs by pruning location. Configuration: $M$ = DinoV2-G on AGX Orin.
-As expected, pruning earlier yields lower latency but greater accuracy degradation.
-For the batch size 1, our method pruned $\sim$54of input tokens at the first layer degraded accuracy by 3.19% but yielded a 40% overall latency reduction.
-Across both batch size 1 and 2 in this ablation study, pruning after the first 25% of layers (layer 10) results in a good balance between latency reduction and accuracy degradation.
-Pruning later in the network will reduce accuracy degradation, however we prioritize yielding latency benefits with our method.
-Therefore, we perform pruning at the layer 25% of the way into the network for all models evaluated in this work.
-That is layer 3 for DeiT-B and DeiT-S, and layer 10 for DinoV2-G.
-Device Batch Model & Acc. Median
-Size Method Loss Latency (ms)
+6 illustrates the total times required for measuring the workload-latency characteristics of DeiT-S and ViT-L across devices.
+The offline computation is relatively fast (no more than 4.5 hours), especially compared with the time required to train any ViT.
+[…]
+Time (min) Time (min)
+TX2 DeiT-S - 54
+ViT-L - 54
+Orin DeiT-S - 54
+ViT-L - 54
+A100 DeiT-S 83 54
+ViT-L 218 52
+[…]
+[…]
+Size Loss Latency (ms)
+TX2 2 DeiT-S -(00.0%) 68.92
+w/ Top-K 4.30 (-27.0%) 50.32
+w/ ToMe 2.47 (-23.1%) 52.98
+w/ DyViT ✝ 0.46 (-26.2%) 50.88
+w/ Ours 1.24 (-28.3%) 49.44
 TX2 2 DeiT-B -(00.0%) 215.0
@@
 w/ Ours 1.16 (-32.1%) 146.0
-TX2 2 DeiT-S -(00.0%) 68.92
-w/ Top-K 4.30 (-27.0%) 50.32
-w/ ToMe 2.47 (-23.1%) 52.98
-w/ DyViT ✝ 0.46 (-26.2%) 50.88
-w/ Ours 1.24 (-28.3%) 49.44
+TX2 2 ViT-L -(00.0%) 1327.0
+w/ Top-K 38.2 (-57.2%) 568.0
+w/ ToMe 17.5 (-57.8%) 559.3
+w/ Ours 8.4 (-55.2%) 594.9
+Orin 4 ViT-L -(00.0%) 70.29
+w/ Top-K 52.70 (-32.2%) 47.63
+w/ ToMe 17.51 (-22.9%) 54.18
+w/ Ours 2.35 (-33.0%) 47.08
 Orin 2 DinoV2-G -(00.0%) 155.5
@@
 w/ Ours 2.04 (-20.1%) 32.37
-Table 6: Comparison with existing methods for similar latency across device and workload sizes.
-We tune the pruning hyperparameters of the compared methods so they achieve similar latency to our method.
-✝DynamicViT (DyViT) requires training.
-Other evaluated methods are training-free.
-4.4 Accuracy-latency tradeoffs of our method
-In this section, we discuss the results of our training-free token pruning mechanism when combined with the pruning schedule inspired by the GPU tail effect.
-We primarily compare against ToMe [1], and include Top-K as a baseline [13, 19].
-DynamicViT [36] is included in select evaluations, however unlike the other compared work and our method it requires training.
+Table 7: Comparison with existing methods for similar latency across devices and workload sizes.
+Hyperparameters of methods were chosen to get as close as possible to match our latency; in some cases pruning more tokens did not reduce latency.
+For each model, the smallest batch size was selected to demonstrate cases where the workload-latency relationship can be exploited.
+✝DynamicViT (DyViT) is the the only training-based method.
+Figure 4: Illustration of accuracy-latency tradeoffs of surveyed methods with $M$ = DinoV2-G: (a) batch size=2 on AGX Orin (b) batch size=4 on A100 (c) batch size=4 on AGX Orin.
+Our pruning schedule and mechanism generate points that expand the pareto front.
+The number of tokens removed at each layer ($r$) of Top-K and ToMe is evaluated from $r=5$ to $=8$ in increments of 1.
+4.3 Comparison to Other Methods
+In this section, we demonstrate the effectiveness of our token pruning schedule and pruning mechanism across devices and workload sizes.
+Our primary focus is on smaller workloads, which are typical in edge deployment scenarios [17].
+For inter-method comparison, we systematically compare to the state-of-the-art method, Token Merging (ToMe) [1].
+We also evaluate two common benchmarks, Top-K [12, 18] and DynamicViT [34].
+We measure Top-K in all conditions; we use DynamicViT only with models for which its pre-trained weights were available.
+DynamicViT is not emphasized as it involves training, making it an unfair comparison with our method, ToMe, and Top-K.
+Accuracy-latency tradeoffs.
+In this section, we discuss the results of our training-free token pruning mechanism and offline computed pruning schedule.
 Tab.
-6 illustrates our method’s ability to retain higher accuracy for similar latency across various devices and workload sizes.
+7 illustrates our method’s ability to retain higher accuracy for similar latency across devices and workload sizes.
 Fig.
-6 demonstrates that our token pruning mechanism and schedule expands the accuracy-latency tradeoffs on the pareto front across various workload sizes and devices.
+4 demonstrates that our token pruning mechanism and schedule expands the accuracy-latency tradeoffs on the pareto front across devices and workload sizes.
 First, our method is able to achieve higher accuracy than other training-free pruning techniques ToMe and Top-K.
@@
 7 is an ablation study across various workload sizes (batch size, models) and devices, where we tune the hyperparameters of methods such that similar latency is achieved.
-Unsurprisingly, DynamicViT is able to retain higher accuracy than all methods, since it requires training.
-Compared with ToMe and Top-K however, it is shown that our method yields consistently lower accuracy degradation.
-Batch size 2 inference on the AGX Orin demonstrates that to achieve similar latency to our method, Top-K degrades accuracy by more than 45.6%, while we degrade accuracy by only 2%.
-Across all workload sizes and devices, ToMe featured 0.47%-5.92% lower accuracy than our method (Tab.
-6).
-Figure 6: Illustration of accuracy-latency tradeoffs of surveyed methods with $M$ = DinoV2-G: (a) batch size=2 on AGX Orin (b) batch size=4 on A100 (c) batch size=4 on AGX Orin.
-Our pruning schedule and mechanism generate points that expand the pareto front.
-The number of tokens removed at each layer ($r$) of Top-K and ToMe is evaluated from $r=5$ to $=8$ in increments of 1.
+Unsurprisingly, DynamicViT retains accuracy since it was finetuned for 300 epochs.
+However, compared to ToMe and Top-K, our method consistently results in lower accuracy degradation.
+In one case, Top-K degrades accuracy by more than 45.6%, while we degrade accuracy by only 2%.
+Across all workload sizes and devices, ToMe had 0.47 to 15.16 lower Top-1 percentage points than our method for similar latency (within 5.2% or 7ms).
 Second, we note that for larger pruning rates such as $r=7$ and $r=8$, ToMe and Top-K remove nearly all input tokens by the last layer of DinoV2-G.
 These high pruning ratios yield high accuracy degradation of over 40% in the case of A100 batch-size 4 for ToMe, as seen in Fig.
-6.
-Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we achieve higher accuracy and lower latency than progressive pruning approaches.
-This shows aggressive pruning at early layers can be strictly better than progressive pruning methods for certain workload sizes.
-4.5 Low workload size observations
-We demonstrate there are some cases where pruning techniques actually increase latency with respect to the baseline model, rather than reduce it.
+4.
+Comparatively, by pruning 54%-75% of tokens at the 10th layer of DinoV2-G according to the tail effect, we achieve higher accuracy and lower latency.
+In both Tab.
+7 and Fig.
+4, our method features lower accuracy degradation than ToMe and Top-K.
+At small workloads, the marginal latency of pruning additional tokens becomes negligible.
+As a result, pruning tokens at each layer degrades accuracy significantly for little latency benefit; existing methods do not account for this behavior.
+Thus, we prune $R$ tokens early; the remaining tokens propagate through the ViT, retaining information which leads to better accuracy.
+Simultaneously, we achieve high latency reduction due to pruning early in the network.
+Low workload size observations.
+For small workloads, token sparsification can actually increase latency due to the overhead associated with token removal mechanisms.
 Tab.
-7 demonstrates an example on the TX2 for single batch inference.
-ToMe increased latency by $\sim$30%, while we reduced latency by $\sim$9%.
-Our tradeoff of 0.99% accuracy degradation and 9% latency reduction is better than ToMe and Top-K, both of which degrade accuracy and provide no latency benefit compared with the baseline DeiT-S pre-trained model.
-For certain workload sizes, using a baseline model or a hardware aware method such as ours is strictly better than attempting to use a token removal method with high overhead.
-Device Method Acc. Median
-Loss Latency (ms)
-TX2 Baseline (-0.00%) 35.32
-Top-K (r=6) 0.73 (+18.6%) 43.38
-ToMe (r=6) 0.27 (+30.3%) 50.70
-Ours (R=69) 0.99 (-9.06%) 32.12
-Table 7: Experiment on DeiT-S with batch size 1 that illustrates pruning overhead.
-Given comparable total pruned tokens (69 to 72), for certain workload sizes a token pruning mechanism may increase latency due to the overhead of performing pruning at multiple layers.
-Our pruning method is applied at a single layer to avoid this relatively significant overhead.
-We found that the operations required to do token pruning add a non-trivial latency overhead for small workload sizes on some devices such as the TX2.
-The overhead from masking or removing tokens can outweigh the latency reduction from processing fewer tokens.
-As explained in Sec.
-3.2, our choice to perform single-layer pruning rather than progressive pruning allows us to yield a latency reduction.
+8 illustrates three examples of this.
+ToMe and Top-K may increase latency by 2-30% with respect to baseline, while we reduce it by 9-26%.
+There are also cases where all methods, including ours, increase latency by 40%-134%.
+We find that for some workloads, using a baseline model or our method is strictly better than attempting to use a token removal method with high overhead.
+[…]
+Size Loss Latency (ms)
+TX2 1 DeiT-S (-0.00%) 35.32
+Top-K 0.73 (+18.6%) 43.38
+ToMe 0.27 (+30.3%) 50.70
+Ours 0.99 (-9.06%) 32.12
+A100 4 ViT-L (-0.00%) 9.49
+Top-K 0.47 (+51.4%) 14.36
+ToMe 0.29 (+134.4%) 22.24
+Ours 0.85 (+40.0%) 13.29
+A100 16 ViT-L -(00.0%) 29.41
+w/ Top-K 0.77 (+2.48%) 30.14
+w/ ToMe 0.51 (+7.65%) 31.66
+w/ Ours 2.26 (-26.3%) 21.68
+Table 8: Experiment that illustrate pruning overhead for certain workload sizes.
+Pruning parameters were chosen such that a similar number of tokens were pruned as our method.
+Token pruning methods may increase latency due to the overhead of pruning itself.
+We reduce overhead through single-layer pruning.
 5 Limitations
-Our approach to token pruning is designed to achieve better tradeoffs when latency-workload is non-linear.
-This occurs primarily for small workload sizes and batch sizes, which are common contexts in edge deployment.
-For large workloads such as DinoV2-G with batch size 256+, our method is less effective because latency-workload relationships becomes linear.
-Last, though we avoid training, our method requires a grid-search.
-In some contexts, this may be expensive (many tokens, large batch size, large model).
-These contexts are outside the focus of this work, but may be relevant for some practitioners.
+Our token pruning approach is optimized for cases where latency-workload relationships are non-linear.
+For large workloads such as DinoV2-G with batch size 256+, our method is less effective because latency-workload relationships becomes linear and more predictable in this case.
 6 Conclusion
-In this work, we investigated how to improve the efficiency of vision transformers by considering latency-workload relationships.
-We made three main contributions.
-First, we provide insight on how ViT latency is affected by underlying hardware, machine learning framework, and kernel launch factors.
-Second, we show how to determine a token pruning schedule by leveraging information about latency-workload relationships.
-Third, we introduced a token pruning mechanism that is training-free and low overhead.
-Our method advances the accuracy-latency Pareto front compared with existing token pruning approaches across workload sizes and devices.
-For single batch inference, other methods may increase latency by 18.6-30.3% with respect to baseline, while we reduce it by 9%.
-For similar latency (within 5.2%) across devices, the state-of-the-art, ToMe, achieves 45.8%-85.4% ImageNet1K accuracy, while we achieve 78.6%-84.5%.
-Improving kernel designs to reduce non-linear latency behavior is a promising future direction.
+In this work, we offered practical guidance on how to improve token pruning for ViTs in the presence of small workloads by utilizing latency-workload relationships.
+We showed how to determine a token pruning schedule by leveraging non-linear latency-workload relationships; in comparison with prior work, our method yields equal or greater latency reductions while maintaining greater accuracy.
+Ultimately, we demonstrated that leveraging workload-latency behavior is effective at improving ViT efficiency via token pruning, especially for small workloads.